### v1

# Magneto-optical transport in tilted type-I double-Weyl semimetals in the presence of orbital magnetic moment

Abstract
The magneto-optical transport of gapless type-I tilted-single Weyl semimetals exhibits suppression of total magnetoconductivities in the presence of orbital magnetic moment in linear and non-linear responses (Yang Gao et al, Phys. Rev. B 105, 165307 (2022)). In this work, we extend our study to investigate magnetoconductivities in gapless type-I tilted double-Weyl semimetals within semiclassical Boltzmann approach and show the differences that arise compared to single Weyl semimetals.
preprint: APS/123-QED
I Introduction
Weyl semimetal is a three-dimensional topological state of matter, in which the conduction and valence bands touch at a finite number of nodes, called Weyl nodes[1, 2, 3, 4]. The Weyl nodes always appear in pairs due to the Nielsen-Ninomiya theorem[5, 5]. Each Weyl node can be regarded as a monopole in k-space carrying the topologcal charge n=1. Weyl semimetal has the Fermi arc surface states that connect the surface projections of two Weyl nodes [6, 7, 8, 9].
However, the topological charge of the Weyl node can be greater than one, namely $J>1$, and the corresponding materials are termed as multi-Weyl semimetals [10, 11, 12, 13]. For J=2, which is refered as double Weyl semimetal (DWSM), the dispersion relation in the vicinity of weyl node is quadratic in two symmetry directions and linear in the third direction. These Weyl nodes are protected by the crystallographic point group symmetries [10]. Density functional theory proposed DWSM in HgCr2Se4 [6] and SrSi2 [11] and can be achieved in photonic crystals [12]. Numerical studies show the presence of multiple surface Fermi arcs in multi-Weyl semimetal [14]. Multi-WSMs show some intriguing transport phenomena [13, 15, 16, 17, 18, 19, 20, 21]
In this paper, we study the linear and nonlinear magneto-optical responses for tilted DWSMs in the presence of an orbital magnetic moment. This has been studied for isotropic or single WSMs [22]. The orbital magnetic moment can be thought as self-rotation of the Bloch wave packet , and modify the energy of the Bloch electron under the external magnetic field [23]. This orbital moment change the magneto-optical responses of tilted Weyl semimetals [22]. We derive an analytic expression for the magnetoconductivity by means of the semiclassical Boltzmann approach. It is found that the orbital magnetic moment induces non-trivial magnetoconductivity term, which gives rise to a partial cancellation of the total magnetoconductivity. This cancellation is more pronounced compared to isotropic WSMs. Further, we analyzed this suppressed feature for linear (B-linear) and quadratic (quadratic-B) contribution in the magnetic field to magnetoconductivities. We also show that the B-linear (quadratic-B) magnetoconductivity exhibits a behavior that is dependent (independent) of the chirality of the Weyl node in both linear or nonlinear response regime as in the case of single WSMs[22].
II Model Hamiltonian and Semiclassical Boltzmann approach
The non-interacting low energy effective Hamiltonian for tilted multi-Weyl semimetals is given by [15, 16, 17, 18, 19, 20, 21],
$\mathcal{H}_{J}$ $=$ $\hbar v_{t}k_{z}\mathsf{1}_{2\times 2}+\alpha_{J}\hbar[(\hat{k}_{-})^{J}\sigma_{+}+(\hat{k}_{+})^{J}\sigma_{-}]+\chi\hbar\mathit{v}_{F}\hat{k}_{z}\sigma_{z}$ (1)
$+$ $\hbar\mathit{v}_{F}t_{s}k_{z}\sigma_{0},$
where $\sigma_{\pm}=\frac{1}{2}(\sigma_{x}\pm i\sigma_{y})$ and $\hat{k}_{\pm}=\hat{k}_{x}\pm i\hat{k}_{y}$, J represents monopole charge, $\mathit{v}_{F}$ is the effective velocity along $\hat{z}$ direction and $\alpha_{J}$ is the material dependent parameter, e.g. $\alpha_{1}$ and $\alpha_{2}$ are the Fermi velocity and inverse of the mass respectively for the isotropic and double WSMs. The energy dispersion for DWSMs is given by $\epsilon_{\bm{k}}^{s}=\hbar t_{s}\mathit{v}_{F}k_{z}+s\hbar\sqrt{\alpha_{n}^{2}(\rho_{2})^{2}+\left(k_{z}\mathit{v}_{F}\right)^{2}}$ with $\rho_{2}=(k_{x}^{2}+k_{y}^{2})$. We will use semiclassical Boltzmann equations in this study.
In the presence of a static magnetic field $\bm{B}$ and a time varying electric field $\bm{E}$, the semiclassical equations of motion are [23, 24]
$\bm{\dot{r}}=\frac{1}{\hbar}\nabla_{\bm{k}}\tilde{\varepsilon}_{\bm{k}}^{s}-\dot{\bm{k}}\times\Omega_{\bm{k}}^{s}$ (2)
$\hbar\bm{\dot{k}}=-e\bm{E}-e\dot{\bm{r}}\times\bm{B}$ (3)
The first term on the right-hand side of Eq. (2) is $\bm{v}_{\bm{k}}^{s}=\frac{1}{\hbar}\bm{\nabla}_{p}\tilde{\varepsilon}_{\bm{k}}^{s}$, defined in terms of an effective band dispersion $\tilde{\varepsilon}_{s}(\bm{k})$. In topological metals such as WSMs, this quantity acquires a term due to the intrinsic orbital moment ,i.e., $\tilde{\epsilon_{\bm{k}}^{s}}=\epsilon_{\bm{k}}^{s}-\bm{m}_{\bm{k}}^{s}\cdot\bm{B}$, while $\bm{m}_{\bm{k}}^{s}$ is the orbital moment induced by the semiclassical “self-rotation” of the Bloch wave packet. The term $\Omega_{\bm{k}}^{s}$ is the Berry curvature [23, 24]
$\bm{\Omega}_{\bm{k}}^{s}$ $=$ $Im[\langle\bm{\nabla}_{k}u_{k}^{s}|\times|\bm{\nabla}_{k}u_{k}^{s}\rangle]$ (4)
$\bm{m}_{\bm{k}}^{s}$ $=$ $-\frac{e}{2\hbar}Im[\langle\bm{\nabla}_{k}u_{k}^{s}|\times(\mathcal{H}_{J}(\bm{k})-\epsilon_{\bm{k}}^{s})|\bm{\nabla}_{k}u_{k}^{s}\rangle]$ (5)
where $|u_{k}^{s}\rangle$ satisfies the equation $\mathcal{H}_{J}(\bm{k})|u_{k}^{s}\rangle=\epsilon_{\bm{k}}^{s}|u_{k}^{s}\rangle$
The general exressions for Berry curvature and orbital magnetic moment for multi-WSMs are [25]
$\bm{\Omega}_{\bm{k}}^{s}=\pm\frac{s}{2}\frac{nv\alpha_{n}^{2}k_{\perp}^{2n-2}}{\beta_{\bm{k},s}^{3}}\{k_{x},k_{y},nk_{z}\}$ (6)
$\bm{m}_{\bm{k},s}^{\pm}=\frac{s}{2}\frac{env\alpha_{n}^{2}k_{\perp}^{2n-2}}{\hbar\beta_{\bm{k},s}^{2}}\{k_{x},k_{y},nk_{z}\}$ (7)
where $\beta_{\bm{k},s}=\sqrt{\alpha_{n}^{2}(\rho_{2})^{2}+\left(k_{z}\mathit{v}_{F}\right)^{2}}$ in the case of DWSMs.
The two equations (2) and (3) can be decoupled to get
$\bm{\dot{r}}=\frac{1}{\hbar D}[\nabla_{\bm{k}}\tilde{\varepsilon}_{\bm{k}}^{s}+e\bm{E}\times\bm{\Omega}_{\bm{k}}^{s})+\frac{e}{\hbar}(\nabla_{\bm{k}}\tilde{\varepsilon}_{\bm{k}}^{s}\cdot\bm{\Omega}_{\bm{k}}^{s})\bm{B}]$ (8)
$\hbar\bm{\dot{k}}=\frac{1}{\hbar D}[-e\bm{E}-\frac{e}{\hbar}\nabla_{\bm{k}}\tilde{\varepsilon}_{\bm{k}}^{s}\times\bm{B}-\frac{e^{2}}{\hbar}(\bm{E}.\bm{B})\bm{\Omega}_{\bm{k}}^{s}]$ (9)
where the factor $D=1+\frac{e}{\hbar}(\bm{\Omega}_{\bm{k}}^{s}\cdot\bm{B})$ modifies the phase space volume.
For a given chirality $s=\pm$ of a single Weyl node, the semiclassical Boltzmann equation (SBE) reads as follows
$\frac{\partial\tilde{f}^{s}}{\partial t}+\bm{\dot{k}}.\frac{\partial\tilde{f}^{s}}{\partial\bm{k}}=\frac{\tilde{f}^{s}-\tilde{f}^{s}_{0}}{\tau}$ (10)
Here, $\tilde{f}^{s}(\tilde{\epsilon_{\bm{k}}^{s}})$ is the electron distribution function. where $\tau$ is the relaxation time originating from the scattering of electrons by phonons, impurities, electrons, and other lattice imperfections [26].
The $\tilde{f}_{0}^{s}(\epsilon_{\bm{k}}^{s})$ can be expanded at low magnetic field as [27]
$\tilde{f}_{0}^{s}(\tilde{\epsilon_{\bm{k}}^{s}})=\tilde{f}_{0}^{s}(\epsilon_{\bm{k}}^{s}-\bm{m}_{\bm{k}}^{s}\cdot\bm{B})$
$\simeq\tilde{f}_{0}^{s}(\epsilon_{\bm{k}}^{s})-\bm{m}_{\bm{k}}^{s}\cdot\bm{B}\frac{\partial\tilde{f}_{0}^{s}(\epsilon_{\bm{k}}^{s})}{\partial\epsilon_{\bm{k}}^{s}}$ (11)
where $\tilde{f}_{0}^{s}(\epsilon_{\bm{k}}^{s})=1/[e^{(\epsilon_{\bm{k}}^{s}-\mu)/k_{B}T}+1]$ with $k_{B}$ the Boltzmann constant, T the temperature, and $\mu$ the chemical potential.
Eq.(10) can be solved by expanding the distribution function as power series in the electric field as
$\tilde{f}^{s}=\tilde{f}_{0}^{s}+\tilde{f}_{1}^{s}e^{-i\omega t}+\tilde{f}_{2}^{s}e^{-2i\omega t}+....$ (12)
where $\tilde{f}_{1}^{s}$ and $\tilde{f}_{2}^{s}$ are the first- and second-order terms for $\bm{E}$, respectively. The electric current density can be calculated by
$\bm{j}=-\frac{e}{(2\pi)^{3}}\int d^{3}kD\bm{\dot{r}}\tilde{f}^{s}$ (13)
III Linear response of double WSMs
For linear electric field response, we retain only first two terms of Eq.(12) and substitute Eq.(3) in Eq.(10)
$\frac{1}{\hbar D}[-e\textbf{E}-\frac{e^{2}}{\hbar}(\textbf{E}\cdot\textbf{B})\Omega_{\bm{k}}^{s}]\cdot\frac{\partial\tilde{f}_{0}^{s}}{\partial\bm{k}}-i\omega\tilde{f}_{1}^{s}=-\frac{\tilde{f}_{1}^{s}}{\tau}$ (14)
Solve for $\tilde{f}_{1}^{s}$, we obtain
$\tilde{f}_{1}^{s}=\frac{\tau}{(1-i\omega\tau)}\frac{1}{\hbar D}[e\textbf{E}+\frac{e^{2}}{\hbar}(\textbf{E}\cdot\textbf{B})\Omega_{\bm{k}}^{s}]\cdot\frac{\partial\tilde{f}_{0}^{s}}{\partial\bm{k}}$ (15)
We expand Eq.(15) up to second order in magnetic field and obtain
$\tilde{f}_{1}^{s}=\frac{\tau}{(1-i\omega\tau)}\biggl{[}e\textbf{E}\cdot\bm{\mathit{v}}_{\bm{k}}^{s}\cdot\frac{\partial\tilde{f}_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}-\frac{e^{2}}{\hbar}(\textbf{B}\cdot\Omega_{\bm{k}}^{s})(\textbf{E}\cdot\bm{\mathit{v}}_{\bm{k}}^{s})\frac{\partial\tilde{f}_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}+\frac{e^{2}}{\hbar}(\textbf{E}\cdot\textbf{B})(\Omega_{\bm{k}}^{s}\cdot\bm{\mathit{v}}_{\bm{k}}^{s})\frac{\partial\tilde{f}_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}-\frac{e}{\hbar}\textbf{E}\cdot\frac{\partial}{\partial\bm{k}}\biggl{(}\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B}\frac{\partial\tilde{f}_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}\biggr{)}$
$-\frac{e^{3}}{\hbar^{2}}(\textbf{B}\cdot\Omega_{\bm{k}}^{s})(\textbf{E}\cdot\textbf{B})(\Omega_{\bm{k}}^{s}\cdot\bm{\mathit{v}}_{\bm{k}}^{s})\frac{\partial\tilde{f}_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}+\frac{e^{3}}{\hbar^{2}}(\textbf{B}\cdot\Omega_{\bm{k}}^{s})^{2}(\textbf{E}\cdot\bm{\mathit{v}}_{\bm{k}}^{s})\frac{\partial\tilde{f}_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}+\frac{e^{2}}{\hbar^{2}}(\textbf{B}\cdot\Omega_{\bm{k}}^{s})\textbf{E}\cdot\frac{\partial}{\partial\bm{k}}\biggl{(}\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B}\frac{\partial\tilde{f}_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}\biggr{)}$
$-\frac{e^{2}}{\hbar^{2}}\Omega_{\bm{k}}^{s}\cdot\frac{\partial}{\partial\bm{k}}\biggl{(}\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B}\frac{\partial\tilde{f}_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}\biggr{)}\biggr{]}$ (16)
From Eqs.(3) and Eq.(III), the expression for current density at time t is given by
$\bm{j}_{1}=-\frac{e}{(2\pi)^{3}}\int d^{3}k\Bigl{[}\bm{\mathit{\tilde{v}}}_{\bm{k}}^{s}+\frac{e}{\hbar}(\Omega_{\bm{k}}^{s}\cdot\bm{\mathit{\tilde{v}}}_{\bm{k}}^{s})\textbf{B}\Bigr{]}\tilde{f}_{1}^{s}$
$-\frac{e^{2}}{2\pi)^{3}\hbar}\int d^{3}k\textbf{E}\times\Omega_{\bm{k}}^{s}\tilde{f}_{0}^{s}$ (17)
The above equation can be expressed in frequency space $\omega$ as
$j_{a}(\omega)=\sigma_{ab}(\omega)E_{b}(\omega)$ (18)
where $\sigma_{ab}(\omega)$ is the frequency dependent conductivity.
III.1 Calculations of longitudinal conductivities components without magnetic field
Substituting Eq.(III) without B terms into the first term of Eq.(III), we get longitudinal conductivities components
$\sigma_{ab}^{(0)}(\omega)=\frac{\tau}{(1-i\omega\tau)}\frac{e^{2}}{(2\pi)^{3}}\int d^{3}k\mathit{v}_{a}^{s}\mathit{v}_{b}^{s}\Bigl{(}-\frac{\partial f_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}\Bigr{)}$ (19)
At T=0 k, $-\frac{\partial f_{0}^{s}}{\partial\varepsilon_{\bm{k}}^{s}}=\delta(\varepsilon_{\bm{k}}^{s}-\mu)$, we get
$\sigma_{zz}^{0}(\omega)$ $=$ $\frac{\hbar\pi\mathit{v}_{F}^{2}\sigma_{D}}{4t_{s}^{2}}\Bigl{[}1-\sqrt{1-t_{s}^{2}}\Bigr{]}$ (20)
$\sigma_{xx}^{0}(\omega)$ $=$ $\sigma_{yy}^{0}(\omega)=\frac{\sigma_{D}}{t_{s}^{3}}\Bigl{[}\frac{t_{s}}{1-t_{s}^{2}}+\frac{1}{2}\ln\frac{1-t_{s}}{1+t_{s}}\Bigr{]}$ (21)
where $\sigma_{D}=\frac{e^{2}\tau\mu^{2}}{(1-i\omega t)2\pi^{2}\mathit{v}_{F}\hbar^{5}}$ is Drude frequency complex conductivity. The $\sigma_{zz}^{0}(\omega)$ are modifed comppared to single WSM while $\sigma_{xx}^{0}(\omega),\sigma_{xx}^{0}(\omega)$ have the same form as in the case of single WSMs.
III.2 Calculations of conductivities components linear in magnetic field
Substituting Eq.(III) with B terms up to first order into the first term of Eq.(III), we get conductivities components
$\sigma_{ab}^{(B)}(\omega)=\sigma_{ab}^{(B,\Omega)}(\omega)+\sigma_{ab}^{(B,m)}(\omega)$ (22)
where
$\sigma_{ab}^{(B,\Omega)}(\omega)=\frac{\tau}{(1-i\omega\tau)}\frac{e^{3}}{(2\pi)^{3}}\int d^{3}k[(\mathit{v}_{a}^{s}B_{b}+\mathit{v}_{b}^{s}B_{a})(\Omega_{\bm{k}}^{s}\cdot\bm{\mathit{v}}_{\bm{k}}^{s})-\mathit{v}_{a}^{s}\mathit{v}_{b}^{s}(\Omega_{\bm{k}}^{s}\cdot\textbf{B})]\Bigl{(}-\frac{\partial f_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}\Bigr{)}$ (23)
$\sigma_{ab}^{(B,m)}(\omega)=\frac{\tau}{(1-i\omega\tau)}\frac{e^{2}}{(2\pi)^{3}}\int d^{3}k\Bigl{[}\frac{\partial\mathit{v}_{a}^{s}}{\partial k_{b}}(\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B})-\frac{\partial(\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B})}{\partial k_{a}}\mathit{v}_{b}^{s}\Bigr{]}\Bigl{(}-\frac{\partial f_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}\Bigr{)}$ (24)
We can easily check from Eqs.(22) and (23) that $\sigma_{ab}^{(B)}(\omega)=\sigma_{ba}^{(B)}(\omega)$. The system possesses the time-reversal symmetry without tilt term and therefore condutivities will vanish [22, 28]. However, for finite value of tilt $t_{s}\neq 0$, the time-reversal symmetry is broken. The first term in Eq. (22) is related to the Berry curvature $\Omega$. The second term in Eq. (22) stems from the contribution of the orbital magnetic moment for isotropic WSMs [22]. In the following, the detailed analysis of Eq. (22) is given by considering the magnetic field B perpendicular and parallel to the tilt directon $t_{s}$.
Case-I When $\textbf{B}\|t_{s}\|z$
In this case, Hall conductivities components are zero and one can get following expressions for longtudinal components
$\sigma_{zz}^{(B,\Omega)}=\frac{\hbar\pi\mathit{v}_{F}^{2}\sigma_{1}^{(B)}s}{3\pi\alpha_{2}\mu}\Bigl{[}\frac{3}{2}\frac{1-2t_{s}^{2}-t_{s}^{4}}{t_{s}^{4}}\ln\frac{1-t_{s}}{1+t_{s}}$
$-\frac{(-3+5t_{s}^{2}-3t_{s}^{3}+3t_{s}^{4})}{t_{s}^{3}}\ln\frac{1-t_{s}}{1+t_{s}}\Bigr{]}$
$\sigma_{xx}^{(B,\Omega)}=\sigma_{1}^{(B)}s\Bigl{[}\frac{(-4+4\sqrt{1-t_{s}^{2}}+t_{s}^{2}(3-\sqrt{1-t_{s}^{2}})}{t_{s}^{5}}\Bigr{]}$
where $\sigma_{1}^{(B)}=\frac{e^{2}\tau\mu eB\alpha_{2}}{\hbar(1-i\omega t)2\pi\mathit{v}_{F}\hbar^{2}}$.
$\sigma_{zz}^{(B,m)}(\omega)$ $=$ $\frac{\hbar\mathit{v}_{F}^{2}\sigma_{0}^{(B)}s}{\pi\alpha_{2}\mu}\Bigl{[}\frac{-3+5t_{s}^{2}}{3t_{s}^{3}}-\frac{(1-t_{s}^{2})^{2}}{2t_{s}^{4}}\ln\frac{1-t_{s}}{1+t_{s}}\Bigr{]}$
$\sigma_{xx}^{(B,m)}(\omega)$ $=$ $\sigma_{0}^{(B)}s\Bigl{[}\frac{(-4-2t_{s}^{4}+4\sqrt{1-t_{s}^{2}}+t_{s}^{2}(7-5\sqrt{1-t_{s}^{2}})}{t_{s}^{5}\sqrt{1-t_{s}^{2}}}\Bigr{]}$ (28)
where $\sigma_{1}^{(B)}=\frac{e^{2}\tau\mu eB\alpha_{2}}{\hbar(1-i\omega t)2\pi\mathit{v}_{F}\hbar^{2}}$.
From the Eqs.(III.2) and (III.2), we have obtained non-trivial dependence of conductivities components on tilt parameter.
Case-II For $\textbf{B}\perp t_{s}\perp z$
We can represent the magnetic field in x-y plane as $\textbf{B}=B(\hat{i}\cos\gamma+\hat{j}\sin\gamma)$ where $\gamma$ is the angle between magnetic field B and x-axis. In this case longitudinal components of condutivities are zero and we get the following expressions for planer Hall conductivities [22, 29]
$\sigma_{xz}^{(B)}(\omega)=[\sigma^{(B,\Omega)}(\omega)+\sigma^{(B,m)}(\omega)]\cos\gamma$ (29)
$\sigma_{yz}^{(B)}(\omega)=[2\sigma^{(B,\Omega)}(\omega)+\sigma^{(B,m)}(\omega)]\sin\gamma$ (30)
$\sigma^{(B,\Omega)}_{xz}(\omega)=\frac{\tau}{(1-i\omega\tau)}\frac{e^{3}}{(2\pi)^{3}}\int d^{3}k[\mathit{v}_{z}^{s}B\cos\gamma(\mathit{v}_{x}^{s}\Omega_{ky}+\mathit{v}_{z}^{s}\Omega_{kz})-\mathit{v}_{z}^{s}B\sin\gamma\mathit{v}_{x}^{s}\Omega_{ky}]\Bigl{(}-\frac{\partial f_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}\Bigr{)}$ (31)
$\sigma^{(B,m)}_{xz}(\omega)=\frac{\tau}{(1-i\omega\tau)}\frac{e^{2}}{(2\pi)^{3}}\int d^{3}k[B\cos\gamma(\frac{\partial\mathit{v}_{x}^{s}}{\partial k_{z}}m_{kx}^{s}-v_{z}^{s}\frac{\partial\mathit{m}_{kx}^{s}}{\partial k_{z}})-B\sin\gamma(\frac{\partial\mathit{v}_{x}^{s}}{\partial k_{z}}m_{ky}^{s}-v_{z}^{s}\frac{\partial\mathit{m}_{ky}^{s}}{\partial k_{x}})]\Bigl{(}-\frac{\partial f_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}\Bigr{)}$ (32)
The second term of both the above equations contribute to zero.
$\sigma^{(B,\Omega)}=\sigma_{0}^{(B)}s\Bigl{[}\frac{-3t_{s}+5t_{s}^{3}-6t_{s}^{5}}{t_{s}^{4}}-\frac{3}{2}\frac{(1-t_{s}^{2})^{2}}{t_{s}^{4}}\ln\frac{1-t_{s}}{1+t_{s}}\Bigr{]}$
$\sigma^{(B,m)}=\sigma_{0}^{(B)}s\Bigl{[}\frac{(3-2t_{s}^{2})}{t_{s}^{3}}+\frac{3}{2}\frac{(1-t_{s}^{2})^{2}}{t_{s}^{4}}\ln\frac{1-t_{s}}{1+t_{s}}\Bigr{]}$
where $\sigma_{0}^{(B)}=\frac{e^{3}\tau B\mathit{v}_{F}}{\hbar(1-i\omega t)12\pi^{2}\hbar}$. From Eqs. (24)–(33), we notice that the B-linear magnetoconductivity in tilted Weyl semimetals is independent of Fermi energy $\mu$, and the odd function of $t_{s}$ as in the case of single WSMs [22]. According to the Nielsen-Ninomiya theorem [5, 5], the Weyl nodes with opposite chirality always appear in pairs. The total magnetoconductivity of the system is the sum of all the Weyl nodes. Therefore, for the case of $t_{+}=t_{-}$ , where the tilt inversion symmetry is broken, the contribution of the Weyl node to the magnetoconductivity has the opposite sign for the opposite (B) chirality , giving rise to $\sigma_{ab}(\omega)$ = 0. Whereas, for the case of $t_{+}=-t_{-}$ , where the tilt inversion symmetry is unbroken,each Weyl node produces an identical contribution to the magnetoconductivity, and so the nonzero magnetoconductivity emerges for this case.
III.3 Calculatons of quadractic-B contributon to the conductivity $\sigma_{ab}^{(B^{2})}$
Substituting Eq.(III) with B terms up to second order into the first term of Eq.(III), we get conductivities components
$\sigma_{ab}^{(B^{2})}(\omega)=\sigma_{ab}^{(B^{2},\Omega)}(\omega)+\sigma_{ab}^{(B^{2},m)}(\omega)$ (34)
where
$\sigma_{ab}^{(B^{2},\Omega)}(\omega)=\frac{\tau}{\hbar^{2}(1-i\omega\tau)}\frac{e^{4}}{(2\pi)^{3}}\int d^{3}k[\mathit{v}_{a}^{s}\mathit{v}_{b}^{s}(\Omega_{\bm{k}}^{s}\cdot\textbf{B})^{2}-(\mathit{v}_{a}^{s}B_{b}+\mathit{v}_{b}^{s}B_{a})(\Omega_{\bm{k}}^{s}\cdot\bm{\mathit{v}}_{\bm{k}}^{s})(\Omega_{\bm{k}}^{s}\cdot\textbf{B})+B_{a}B_{b}(\Omega_{\bm{k}}^{s}\cdot\bm{\mathit{v}}_{\bm{k}}^{s})^{2}]\Bigl{(}-\frac{\partial f_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}\Bigr{)}$
$\sigma_{ab}^{(B^{2},m)}(\omega)=\frac{\tau}{\hbar^{2}(1-i\omega\tau)}\frac{e^{3}}{(2\pi)^{3}}\int d^{3}k\Bigl{[}\frac{\partial}{\partial\textbf{k}}\cdot[\mathit{v}_{a}^{s}B_{b}\Omega_{\bm{k}}^{s}](\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B})-\frac{\partial[\mathit{v}_{a}^{s}(\Omega_{\bm{k}}^{s}\cdot\textbf{B})]}{\partial k_{b}}(\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B})+\frac{\partial[B_{a}^{s}(\Omega_{\bm{k}}^{s}\cdot\mathit{\bm{v}}_{\bm{k}}^{s})]}{\partial k_{b}}(\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B})$
$+\frac{\partial(\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B})}{\partial k_{a}}(\Omega_{\bm{k}}^{s}\cdot\textbf{B})\mathit{v}_{b}^{s}-\frac{\partial(\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B})}{\partial k_{a}}(\Omega_{\bm{k}}^{s}\cdot\mathit{\bm{v}}_{\bm{k}}^{s})B_{b}^{s}-\frac{\partial^{2}(\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B})}{e\partial k_{a}\partial k_{b}}(\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B})-B_{a}\frac{\partial(\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B})}{\partial\bm{k}}\cdot\Omega_{\bm{k}}^{s}\mathit{v}_{b}^{s}\Bigr{]}\Bigl{(}-\frac{\partial f_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}\Bigr{)}$
$\sigma_{ab}^{(B^{2},\Omega)}(\omega)$ and $\sigma_{ab}^{(B^{2},m)}(\omega)$ Berry-curvature $\Omega_{\bm{k}}^{s}$ and the orbital magnetic moment $\textbf{m}_{\bm{k}}^{s}$ dependent conductivities respectively. Consider the fllowing two cases
Case-I When $\textbf{B}\|t_{s}\|z$
In this case, Hall conductivities components are zero and one can get following expressions for longtudinal components
$\sigma_{zz}^{(B^{2},\Omega)}(\omega)=\frac{5\pi\mathit{v}_{F}^{2}\hbar}{4\mu\alpha_{2}}\sigma_{0}^{(B^{2})}$
$\sigma_{xx}^{(B^{2},\Omega)}(\omega)=8\sigma_{1}^{(B^{2})}\Bigl{[}\frac{-2(15-25t_{s}^{2}+8t_{s}^{4})}{t_{s}^{6}}$
$-15\frac{(1-t_{s}^{2})^{2}}{t_{s}^{7}}\ln\frac{1-t_{s}}{1+t_{s}}\Bigr{]}$ (37)
Considering the effect of the orbital magnetic moment, we have
$\sigma_{xx}^{(B^{2},m)}(\omega)=-\frac{\pi\mathit{v}_{F}^{2}\hbar}{\mu\alpha_{2}}\sigma_{0}^{(B^{2})}$ (38)
$\sigma_{zz}^{(B^{2},m)}(\omega)=\frac{16}{5}\sigma_{1}^{(B^{2})}\Bigl{[}\frac{(30-20t_{s}^{2}+t_{s}^{4})}{t_{s}^{6}}$
$+\frac{5}{2}\frac{(6-t_{s}^{2}+t_{s}^{4})}{t_{s}^{7}}\ln\frac{1-t_{s}}{1+t_{s}}]$ (39)
where $\sigma_{1}^{(B^{2})}=\frac{e^{2}\tau(eB)^{2}\alpha^{2}}{8\pi^{2}\hbar^{2}(1-i\omega t)\hbar\mathit{v}_{F}}$.
Case-II In this case both longitudinal components as well as planer Hall conductivities are non-zero. The longtudnal components have the following expressions
$\sigma_{xx}^{(B^{2},\Omega)}(\omega)$ $=$ $\frac{\pi\mathit{v}_{F}^{2}\hbar}{32\mu\alpha_{2}}\sigma_{1}^{(B^{2})}[(31+16t_{s}^{2})\cos^{2}\gamma+5\pi^{2}\sin^{2}\gamma]$
$\sigma_{zz}^{(B^{2},\Omega)}(\omega)$ $=$ $\sigma_{1}^{(B^{2})}\frac{\mathit{v}_{F}^{4}\hbar^{2}}{15\mu^{2}}(1+7t_{s}^{2})$
The off-diagonal components of conductivities
$\sigma_{xy}^{(B^{2},\Omega)}(\omega)=\sigma_{1}^{(B^{2})}\frac{\pi\mathit{v}_{F}^{2}\hbar}{16\mu\alpha_{2}}(13+8t_{s}^{2})\sin\gamma\cos\gamma$ (42)
The magnetic orbital moment contrbutions of conductivities
$\sigma_{xx}^{(B^{2},m)}(\omega)=\frac{\pi\mathit{v}_{F}^{2}}{2\mu\alpha_{2}}\sigma_{0}^{(B^{2})}(2\cos^{2}\gamma-\sin^{2}\gamma)$ (43)
$\sigma_{zz}^{(B^{2},m)}(\omega)=-\frac{\mathit{v}_{F}^{4}\hbar^{2}}{15\mu^{2}\alpha_{2}^{2}}\sigma_{1}^{(B^{2})}(1-t_{s}^{2})$ (44)
$\sigma_{xy}^{(B^{2},m)}(\omega)=\frac{2\pi^{2}\mathit{v}_{F}^{2}\hbar}{\mu\alpha_{2}}\sin\gamma\cos\gamma\sigma_{1}^{(B^{2})}$ (45)
It is noted that all the other magnetoconductivity components are zero. It is clear that the above conductivity equations are independent of chirality, i.e., the Weyl cones with opposite chiralities have the same contribution to the conductivity. In contrast to single WSMs, only the conductivity component $\sigma_{zz}^{(B^{2},m)}(\omega)$ is always negative, a result that will leave the total conductivity weakened. Other components of conducivities are enhanced due to orbital magnetic moment.
III.4 Hall conductivities $\sigma_{ab}^{(H,0)}$ and $\sigma_{ab}^{(H,B)}$
The intrinsic Hall effect is engendered by the Berry curvature, as presented by the second term of Eq. (III), for which the further calculation gives
$\sigma_{ab}^{(H)}=\sigma_{ab}^{(H,0)}+\sigma_{ab}^{(H,B)}$ (46)
where
$\sigma_{ab}^{(H,0)}(\omega)$ $=$ $-\frac{e^{2}}{\hbar(2\pi)^{3}}\epsilon_{abc}\int d^{3}k\Omega_{c}^{s}f_{0}^{s}$ (47)
$\sigma_{ab}^{(H,0)}(\omega)$ $=$ $\frac{e^{2}}{\hbar(2\pi)^{3}}\epsilon_{abc}\int d^{3}k\Omega_{c}^{s}(\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B})\Bigl{(}-\frac{\partial f_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}\Bigr{)}$ (48)
where $\epsilon_{abc}$ is the Levi-Civita symbol with $a,b,c,\in{x,y,z}$ The first term $\sigma_{ab}^{(H,0)}(\omega)$ in Eq. (46), referring to the anomalous Hall effect, is not equivalent to zero only in the system with broken time-reversal symmetry [22]. While the second term $\sigma_{ab}^{(H,0)}(\omega)$stands for the ordinary Hall conductivity linear in B, which is counterpart to a semiclassical description related to Landau level formation in the quantum limit.
Let the magnetic field in spherical polar co-ordinate system $\textbf{B}=(B_{x},B_{y},B_{z})$ with $B_{x}=B\sin\theta\cos\phi,B_{x}=B\sin\theta\sin\phi,B_{z}=B\cos\theta$. For a single Weyl node, the B-linear contribution to the Hall conductivity can be written as
$\sigma^{(H,B)}\\ =\left(\begin{array}[]{ccc}0&\sigma_{1}^{H}\cos\theta&-\sigma_{1}^{H}\sin\theta\sin\phi\\ -\sigma_{1}^{H}\cos\theta&0&\sigma_{1}^{H}\sin\theta\cos\phi\\ \sigma_{1}^{H}\sin\theta\sin\phi&-\sigma_{1}^{H}\sin\theta\cos\phi&0\end{array}\right)$ (49)
where $\sigma_{1}^{H}=-\frac{e^{2}}{\hbar}\frac{Be}{4\pi}\frac{(-2+t_{s}^{2}+\sqrt{1-t_{s}^{2}})}{t_{s}^{4}\mathit{v}_{F}\hbar}\alpha_{2}$
In contrast to the single WSM node, the B-linear contribution to the Hall conductivity of Weyl semimetals depends on the tilt and depends on the chirality of the valley as in the case of single WSM. The linear Hall conductivity is independent of $\mu$ while in case of single WSM, it is inversely proportional to $\mu$.
IV Second order non-linear response of DWSMs
Now, we exlore the second-order nonlinear magneto-optical response of Weyl semimetals. Substituting Eq.(9) into Eq.(10), neglecting the Lorentz force term, and retaining terms up to second order in E, we obtain
$\frac{1}{\hbar D}[-e\textbf{E}-\frac{e^{2}}{\hbar}(\textbf{E}\cdot\textbf{B})\Omega_{\bm{k}}^{s}]\cdot\frac{\partial\tilde{f}_{1}^{s}}{\partial\bm{k}}-i\omega\tilde{f}_{2}^{s}=-\frac{\tilde{f}_{2}^{s}}{\tau}$ (50)
Solve for $\tilde{f}_{2}^{s}$, we obtain
$\tilde{f}_{2}^{s}=\frac{\tau}{(1-2i\omega\tau)}\frac{1}{\hbar D}[e\textbf{E}+\frac{e^{2}}{\hbar}(\textbf{E}\cdot\textbf{B})\Omega_{\bm{k}}^{s}]\cdot\frac{\partial\tilde{f}_{1}^{s}}{\partial\bm{k}}$ (51)
Substituting Eq.(15) into Eq.(51) and retan terms up to first order in B, we obtain
$\tilde{f}_{2}^{s}=\frac{e\tau^{2}}{(1-i\omega\tau)(1-2i\omega\tau)}\biggl{\{}\textbf{E}\cdot\frac{\partial}{\partial\bm{k}}\biggr{[}\Bigl{(}e\textbf{E}+\frac{e^{2}}{\hbar}(\textbf{E}\cdot\textbf{B})\Omega_{\bm{k}}^{s}-\frac{e^{2}}{\hbar}(\textbf{B}\cdot\Omega_{\bm{k}}^{s})\textbf{E}\Bigr{)}\cdot\bm{\mathit{v}}_{\bm{k}}^{s}\frac{\partial f_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}-\frac{e}{\hbar}\textbf{E}\cdot\frac{\partial}{\partial\bm{k}}\biggl{(}\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B}\frac{\partial f_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}\biggr{)}\biggr{]}$
$+\biggl{[}\frac{e^{2}}{\hbar}(\textbf{E}\cdot\textbf{B})\Omega_{\bm{k}}^{s}-\frac{e^{2}}{\hbar}(\textbf{B}\cdot\Omega_{\bm{k}}^{s})\textbf{E}\biggr{]}\cdot\frac{\partial}{\partial\bm{k}}\biggl{(}\textbf{E}\cdot\bm{\mathit{v}}_{\bm{k}}^{s}\frac{\partial f_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}\biggr{)}\biggr{\}}$ (52)
Now the electric current density at the frequency $2\omega$is given by
$\bm{j}_{1}=-\frac{e}{(2\pi)^{3}}\int d^{3}k\Bigl{[}\bm{\mathit{\tilde{v}}}_{\bm{k}}^{s}+\frac{e}{\hbar}(\Omega_{\bm{k}}^{s}\cdot\bm{\mathit{\tilde{v}}}_{\bm{k}}^{s})\textbf{B}\Bigr{]}\tilde{f}_{2}^{s}$
$-\frac{e^{2}}{2\pi)^{3}\hbar}\int d^{3}k\textbf{E}\times\Omega_{\bm{k}}^{s}\tilde{f}_{1}^{s}$ (53)
According to the definition of second harmonic conductivity, this equation should be written in the form
$\textbf{j}(2\omega)=\sigma(2\omega)\textbf{E}(\omega)\textbf{E}(\omega)$ (54)
where $\sigma(2\omega)$ is the second harmonic conductivity.
IV.1 second harmonic conductivity $\sigma_{abc}^{0}$ in the absence of magnetic field
In this subsection, we calculate the second harmonic current of the Weyl semimetals system in absence of magnetic fields $\textbf{B}=0$. Inserting Eq. (IV) into the first term of Eq. (IV), the second harmonic conductivity tensor can be written as
$\sigma_{abc}^{0}(2\omega)=\frac{-\tau^{2}}{(1-i\omega\tau)(1-2i\omega\tau)}\frac{e^{3}}{\hbar(2\pi)^{3}}\int d^{3}k\frac{\partial\mathit{v}_{a}^{s}}{\partial k_{c}}\mathit{v}_{b}^{s}\Bigl{(}-\frac{\partial f_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}\Bigr{)}$ (55)
Except the aaz, aza, and zaa (a = x, y, z) components of the second harmonic conductivity tensor are nonzero and all other components equal to zero [22, 30]
$\sigma_{zzz}^{0}(2\omega)=-\frac{3\pi\mathit{v}_{F}^{2}}{8\alpha_{2}}\sigma_{DL}\Bigl{[}\frac{2-2\sqrt{1-t_{s}^{2}}}{t_{s}^{3}}+\frac{-3+2\sqrt{1-t_{s}^{2}}}{t_{s}}\Bigr{]}$ (56)
$\sigma_{xxz}^{0}(2\omega)=\sigma_{DL}\Bigl{[}\frac{3-2t_{s}^{2}}{t_{s}^{3}}+3\frac{1-t_{s}^{2}}{2t_{s}^{4}}\ln\frac{1-t_{s}}{1+t_{s}}\Bigr{]}$ (57)
where $\sigma_{DL}=\frac{e^{3}\tau^{2}}{\hbar(1-2\omega t)(1-i\omega t)6\pi^{2}\hbar}$ is the Drude-like frequency dependent complex conductivity.
The second harmonic conductivity tensor satisfies the relation $\sigma_{xxz}^{0}(2\omega)\sigma_{yyz}^{0}(2\omega)=\sigma_{zxx}^{0}(2\omega)=\sigma_{zyy}^{0}(2\omega)=\sigma_{xzx}^{0}(2\omega)=\sigma_{yzy}^{0}(2\omega)$,$\sigma_{xxx}^{0}(2\omega)=\sigma_{yyy}^{0}(2\omega)=0$ and it does not depend on the chirality of Weyl node. The total second harmonic conductivity in tilted Weyl semimetals is the sum of a pair of Weyl nodes. For the case with tilt inversion symmetry $t_{+}=-t_{-},\sigma_{abc}(2\omega)=0$. For the case with broken tilt inversion symmetry $t_{+}=t_{-},\sigma_{abc}(2\omega)\neq 0$.
IV.2 second harmonic conductivity $\sigma_{abc}^{B}$ in the presence of linear magnetic field
Inserting Eq. (IV) into the first term of Eq. (IV), we write the lnear B dependent second harmonic conductivity tensor
$\sigma_{abc}^{B}(2\omega)=\frac{\tau^{2}}{(1-i\omega\tau)(1-2i\omega\tau)}\frac{e^{3}}{\hbar(2\pi)^{3}}\int d^{3}k\biggl{\{}\frac{\partial\mathit{v}_{a}^{s}}{\partial k_{c}}\Bigl{[}-\frac{eB_{b}}{\hbar}(\Omega_{\bm{k}}^{s}\cdot\bm{\mathit{v}}_{\bm{k}}^{s})+\frac{e\mathit{v}_{b}^{s}}{\hbar}(\Omega_{\bm{k}}^{s}\cdot\bm{B})\Bigr{]}-\frac{eB_{c}}{\hbar}\frac{\partial}{\partial\bm{k}}\cdot(\mathit{v}_{a}^{s}\Omega_{\bm{k}}^{s})\mathit{v}_{b}^{s}$
$+\frac{e}{\hbar}\frac{\partial[\mathit{v}_{a}^{s}(\Omega_{\bm{k}}^{s}\cdot\bm{B})]}{\partial k_{c}}\mathit{v}_{b}^{s}-\frac{eB_{a}}{\hbar}\frac{\partial(\Omega_{\bm{k}}^{s}\cdot\bm{\mathit{v}}_{\bm{k}}^{s})}{\partial k_{c}}\mathit{v}_{b}^{s}+\frac{\partial^{2}(\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B})}{\hbar\partial k_{a}\partial k_{c}}\mathit{v}_{b}^{s}-\frac{\partial^{2}\mathit{v}_{a}^{s}}{\hbar\partial k_{a}\partial k_{c}}(\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B})\biggr{\}}\Bigl{(}-\frac{\partial f_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}\Bigr{)}$ (58)
We will explore Eq.(IV.2) with follwing two cases:
Case-I In the case of $\textbf{B}\parallel t_{s}$ , for a single Weyl node, we obtain the magnetoconductivity components
$\sigma_{xzx}^{B}(2\omega)=2s\sigma_{2}^{(B)}\frac{\sqrt{1-t_{s}^{2}}}{t_{s}^{4}}[2-t_{s}^{2}-2\sqrt{1-t_{s}^{2}}]$ (59)
$\sigma_{zxx}^{B}(2\omega)=-s\sigma_{2}^{(B)}\frac{1}{t_{s}^{4}}[2-t_{s}^{2}-2\sqrt{1-t_{s}^{2}}]$ (60)
where $\sigma_{2}^{(B)}=\frac{e^{3}\tau^{2}\alpha_{2}}{\hbar(1-2\omega t)(1-i\omega t)}\frac{eB}{4\pi\hbar}$
We can check that other nonzero second harmonic conductivity components satisfy the relations: $\sigma_{zxx}^{B}(2\omega)=\sigma_{zyy}^{B}(2\omega)=\sigma_{xzx}^{B}(2\omega)=\sigma_{yzy}^{B}(2\omega)$. In contrast to single WSM, the B-linear contribution to the second harmonic conductivity deends on the tilt. Summing the conductivity over the Weyl cones with opposite chirality cancels each other, leading to the disappearance of the total B-linear contribution to the second harmonic conductivity.
Case-II In the case of $\textbf{B}\perp t_{s}$ , we get the nonzero conductivity components
$\sigma_{zxz}^{(B)}(2\omega)=s\frac{2\mathit{v}_{F}^{2}}{3\pi\mu\alpha_{2}}\sigma_{2}^{(B)}\cos\gamma$ (61)
$\sigma_{xzz}^{(B)}(2\omega)=-s\frac{4\mathit{v}_{F}^{2}}{3\pi\mu\alpha_{2}}\sigma_{2}^{(B)}\cos\gamma$ (62)
$\sigma_{zyz}^{(B)}(2\omega)=s\frac{2\mathit{v}_{F}^{2}}{3\pi\mu\alpha_{2}}\sigma_{2}^{(B)}\sin\gamma$ (63)
$\sigma_{yzz}^{(B)}(2\omega)=-s\frac{8\mathit{v}_{F}^{2}}{3\pi\mu\alpha_{2}}\sigma_{2}^{(B)}\sin\gamma$ (64)
In this case, the B-linear contribution to the second harmonic conductivity is dependent of the chirality but independent of the tilt. Summing the conductivity over the Weyl cones with opposite chirality cancels each other, leading to the disappearance of the total B-linear contribution to the second harmonic conductivity.
IV.3 The second order nonlinear Hall conductivity $\sigma_{abc}^{(H,0)}$ in the absence of magnetc field
In this subsection, we study the second-order nonlinear Hall effect of Weyl semimetals wthout magnetic field. Inserting Eq. (III) into the second term in Eq. (IV) and taking B = 0, we obtain the nonlinear Hall conductivity.
$\sigma_{abc}^{(H,0)}=\epsilon_{acd}\frac{e^{3}\tau}{\hbar^{2}(1-i\omega\tau)}D_{bd}$ (65)
where $D_{bd}$ is called Berry dipole. It is a first order moment of the Berry curvature[31].
One can calculate the non-zero components of Berry curvature dipole
$D_{bd}=\frac{\hbar}{(2\pi)^{3}}\int d^{3}k\Omega_{d}^{s}\mathit{v}_{b}^{s}\Bigl{(}-\frac{\partial f_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}\Bigr{)}$ (66)
$D_{xx}=D_{yy}=-s\frac{1}{8\pi^{2}}\Bigl{[}\frac{2}{t_{s}^{2}}+\frac{1-t_{s}^{2}}{t_{s}^{3}}\ln\frac{1-t_{s}}{1+t_{s}}\Bigr{]}$ (67)
$D_{zz}=-s\frac{1}{4\pi^{2}}\frac{t_{s}^{2}-1}{t_{s}^{3}}\Bigl{[}2t_{s}+\ln\frac{1-t_{s}}{1+t_{s}}\Bigr{]}$ (68)
In the case of DWSMs,the magntude of Berry curvature dipole components are twice the values of single WSMs and it has the same dependence on the chirality, as well as on the tilt. Therefore, contributions from a pair of Weyl nodes with opposite chirality exactly cancel each other [22, 31, 32].
IV.4 Linear B-contribution to the second order nonlinear Hall conductivity $\sigma_{abc}^{(H,B)}$
Now, we focus on the second-order nonlinear Hall effect in a weak magnetic field. Inserting Eq. (III) into the second term of Eq. (IV) and retain terms up to first power of B, one obtan complex nonlinear Hall conductivity
$\sigma_{abc}^{(H,B)}=\epsilon_{acd}\frac{e^{3}\tau}{\hbar^{2}(1-i\omega\tau)}\bigl{[}D_{bd}^{\Omega}+D_{bd}^{m}\bigr{]}$ (69)
where
$D_{bd}^{\Omega}=\frac{e}{(2\pi)^{3}}\int d^{3}k\Omega_{d}^{s}[B_{b}(\Omega_{\bm{k}}^{s}\cdot\bm{\mathit{v}}_{\bm{k}}^{s})-\mathit{v}_{b}^{s}(\Omega_{\bm{k}}^{s}\cdot\bm{B})]\Bigl{(}-\frac{\partial f_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}\Bigr{)}$ (70)
$D_{bd}^{m}=\frac{1}{(2\pi)^{3}}\int d^{3}k\frac{\partial\Omega_{d}^{s}}{\partial k_{b}}(\bm{m}_{\bm{k}}^{s}\cdot\bm{B})\Bigl{(}-\frac{\partial f_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}\Bigr{)}$ (71)
where $D_{bd}^{\Omega}$ and $D_{bd}^{m}$ are Berry curvature dipole contributions due to the Berry-curvature and the orbital magnetic moment respectively.
Again we will consider the following two cases.
Case-I In the case of $\textbf{B}\parallel t_{s}$ , one obtain the Berry curvature dipole components
$D_{xx}^{\Omega}=D_{yy}^{\Omega}=D_{zz}^{\Omega}=0$ (72)
$D_{xx}^{m}=D_{yy}^{m}=D_{zz}^{m}=0$ (73)
In this case all the components are zero which are finite in case of isotropic WSM[22].
Case-II In the case of $\textbf{B}\perp t_{s}$ , we get the nonzero components
$D_{zx}^{\Omega}$ $=$ $-\frac{4}{3}t_{s}D_{2}^{(B)}\cos\gamma$ (74)
$D_{xz}^{\Omega}$ $=$ $\frac{15\pi}{2}\frac{\mu\alpha_{2}}{\mathit{v}_{F}^{2}\hbar}t_{s}D_{2}^{(B)}\cos\gamma$ (75)
$D_{zy}^{\Omega}$ $=$ $-\frac{4}{3}t_{s}D_{2}^{(B)}\sin\gamma$ (76)
$D_{yz}^{\Omega}$ $=$ $-\frac{15\pi}{2}\frac{\mu\alpha_{2}}{\mathit{v}_{F}^{2}\hbar}t_{s}D_{2}^{(B)}\sin\gamma$ (77)
and
$D_{zx}^{m}=-6t_{s}D_{2}^{(B)}\cos\gamma$ (78)
$D_{zy}^{m}=-6t_{s}D_{2}^{(B)}\sin\gamma$ (79)
where $D_{2}^{(B)}=\frac{1}{8\pi^{2}}\frac{eB}{\hbar}\frac{\hbar^{2}\mathit{v}_{F}^{2}}{15\mu^{2}}$. One can check that $D_{zx}^{m}=D_{xz}^{m}$ and $D_{zy}^{m}=D_{yz}^{m}$ and rest of the components will vanish. The contributions of all components are renormalized due to parameters $\alpha_{2}$ and $\mathit{v}_{F}$ of DWSMs as compared to single WSMs [22].
The nonlinear Hall effect can be modulated by the polarization of the incident light, as discussed in Ref.[22]. Using Eq. (69),the electric current is rewritten in the form of
$\textbf{j}(2\omega)=\frac{e^{3}\tau}{\hbar^{2}(1-i\omega\tau)}(\hat{D}\cdot\textbf{E})\times\textbf{E}$ (80)
Assume that an electromagnetic wave propogates in the x direction:
$\textbf{E}(\textbf{r},t)=|\textbf{E}(\omega)|Re[|\psi\rangle e^{i(qx-\omega t)}$ (81)
where
$|\psi\rangle=\Bigl{(}\begin{array}[]{c}\psi_{y}\\ \psi_{z}\end{array}\Bigr{)}=\Bigl{(}\begin{array}[]{c}\sin\theta e^{i\alpha_{y}}\\ \cos\theta e^{i\alpha_{z}}\end{array}\Bigr{)}$ (82)
is the Jones vector in the y-z plane with phases $\alpha_{y},\alpha_{z}$ , and the amplitudes $E_{y}=|\textbf{E}|\sin\theta$ and $E_{z}=|\textbf{E}|\cos\theta$ Inserting Eq. (81) into Eq. (80), we obtain the nonlinear Hall current
$j_{x}(2\omega)=\frac{e^{3}\tau}{\hbar^{2}(1-i\omega\tau)}\frac{D_{yy}^{(B)}-D_{zz}^{(B)}}{2}\sin 2\theta e^{(\alpha_{y}+\alpha_{z})}|\textbf{E}|^{2}$ (83)
Following the same procedure of Ref.([22]), Eq.(83) can be rewritten as $j_{x}\sim(D_{yy}-D_{zz})E_{y}E_{z}$ Hence conductivity will be
$\sigma_{x}^{(H,B)}(2\omega)=0$ (84)
In contrast to the case of single WSM, we get zero value of conductivity.
V Conclusion
We have calculated the magneto-optical conductivities of gapless type-I tilted-double Weyl semimetals in the presence of orbital magnetic moment in linear and non-linear responses within semiclassical Boltzmann approach and compared it with case of single WSMs. We have observed that condutivites components are renormalzed in non-trivial manner as compared to single WSMs. Our calculations would be tested in DWSMs materials.