### v2

# Magneto-optical transport in tilted type-I multi-Weyl semimetals in the presence of orbital magnetic moment

Abstract
The magneto-optical transport of gapless type-I tilted single Weyl semimetals exhibits suppression of total magnetoconductivities in the presence of orbital magnetic moment(OMM) in linear and nonlinear responses (Yang Gao et al., Phys. Rev. B 105, 165307 (2022)). In this work, we extend our study to investigate magnetoconductivities in gapless type-I tilted multi-Weyl semimetals(mWSMs) within the semiclassical Boltzmann approach and show the differences that arise compared to single-Weyl semimetals.
preprint: APS/123-QED
I Introduction
Weyl semimetal is a three-dimensional topological state of matter, in which the conduction and valence bands touch at a finite number of nodes, called Weyl nodes[1, 2, 3, 4]. The Weyl nodes always appear in pairs due to the Nielsen-Ninomiya theorem[5, 5]. Each Weyl node can be regarded as a monopole in k-space carrying the topological charge n=1. Weyl semimetal has the Fermi arc surface states that connect the surface projections of two Weyl nodes [6, 7, 8, 9].
However, the topological charge of the Weyl node can be greater than one, namely $J>1$, and the corresponding materials are termed as multi-Weyl semimetals(mWSMs) [10, 11, 12, 13]. For J=2, which is referred to as double Weyl semimetal (DWSM), the dispersion relation in the vicinity of the Weyl node is quadratic in two symmetry directions and linear in the third direction. These Weyl nodes are protected by the crystallographic point group symmetries [10]. Density functional theory proposed DWSM in HgCr2Se4 [6] and SrSi2 [11] and can be achieved in photonic crystals [12]. Numerical studies show the presence of multiple surface Fermi arcs in multi-Weyl semimetal [14]. Multi-WSMs show some intriguing transport phenomena [13, 15, 16, 17, 18, 19, 20, 21]
In this paper, we study the linear and nonlinear magneto-optical responses for tilted mWSMs in the presence of an orbital magnetic moment. This has been studied for isotropic or single WSMs [22]. However, non-linear magnetoconductivity has not been discussed in the literature with the combined effects of both the tilting and orbital magnetic moment terms [23, 24, 25]. The orbital magnetic moment can be thought of as the self-rotation of the Bloch wave packet, and modifies the energy of the Bloch electron under the external magnetic field [26]. This orbital moment changes the magneto-optical responses of tilted mWSMs [22]. We derive an analytic expression for the magnetoconductivity employing the semiclassical Boltzmann approach. It is found that the orbital magnetic moment induces a non-trivial magnetoconductivity term, which gives rise to a partial cancellation of the total magnetoconductivity. This cancellation is more pronounced compared to isotropic WSMs. Further, we analyzed this suppressed feature for linear and quadratic contributions in the magnetic field to magnetoconductivities. We also show that the linear-B (quadratic-B) magnetoconductivity exhibits a behavior that is dependent (independent) of the chirality of the Weyl node in both linear or nonlinear response regimes, as in the case of single WSMs[22].
The paper is organized as follows: In Sec.II, we begin with the model of a 3D multi Weyl semimetal with a tilt in the z direction, and then the semiclassical equations of motion for the dynamics of the electron wave packet in the electric and magnetic fields are presented. In Sec. III, the B-linear and quadratic-B magnetoconductivities including the orbital magnetic moment are obtained in the linear response regime, and analyzed in detail. In Sec. IV, we study second harmonic generation, and give the second harmonic conductivity formula as well as the further analysis for this result. We end with conclusions in Sec. V.
II Model Hamiltonian and Semiclassical Boltzmann approach
The non-interacting low-energy effective Hamiltonian for tilted multi-Weyl semimetals is given by [15, 16, 17, 18, 19, 20, 21],
$\mathcal{H}_{J}=\alpha_{J}\hbar[(k_{-})^{J}\sigma_{+}+(k_{+})^{J}\sigma_{-}]+\chi\hbar\mathit{v}_{F}k_{z}\sigma_{z}+\hbar\mathit{v}_{F}t_{s}k_{z}\sigma_{0}$
where $\sigma_{\pm}=\frac{1}{2}(\sigma_{x}\pm i\sigma_{y})$ and $k_{\pm}=k_{x}\pm ik_{y}$, J represents monopole charge, $\mathit{v}_{F}$ is the effective velocity along $\hat{z}$ direction and $\alpha_{J}$ is the material dependent parameter, e.g. $\alpha_{1}$ and $\alpha_{2}$ are the Fermi velocity and inverse of the mass respectively for the isotropic and double WSMs. The energy dispersion for mWSMs is given by $\epsilon_{\bm{k}}^{s}=\hbar t_{s}\mathit{v}_{F}k_{z}+s\hbar\sqrt{\alpha_{J}^{2}k_{\perp}^{2J}+\left(k_{z}\mathit{v}_{F}\right)^{2}}$ with $k_{\perp}=\sqrt{k_{x}^{2}+k_{y}^{2}}$. We will use semiclassical Boltzmann equations in this study.
In the presence of a static magnetic field $\bm{B}$ and a time varying electric field $\bm{E}$, the semiclassical equations of motion at the location r and the wave-vector k in a given band are [26, 27]
$\bm{\dot{r}}=\frac{1}{\hbar}\nabla_{\bm{k}}\tilde{\varepsilon}_{\bm{k}}^{s}-\dot{\bm{k}}\times\Omega_{\bm{k}}^{s}$ (2)
$\hbar\bm{\dot{k}}=-e\bm{E}-e\dot{\bm{r}}\times\bm{B}$ (3)
where -e is the electron charge. The first term on the right-hand side of Eq. (2) is $\bm{v}_{\bm{k}}^{s}=\frac{1}{\hbar}\bm{\nabla}_{p}\tilde{\varepsilon}_{\bm{k}}^{s}$, defined in terms of an effective band dispersion $\tilde{\varepsilon}_{s}(\bm{k})$. In topological metals such as WSMs, this quantity acquires a term due to the intrinsic orbital moment,i.e., $\tilde{\epsilon_{\bm{k}}^{s}}=\epsilon_{\bm{k}}^{s}-\bm{m}_{\bm{k}}^{s}\cdot\bm{B}$, while $\bm{m}_{\bm{k}}^{s}$ is the orbital moment induced by the semiclassical “self-rotation” of the Bloch wave packet. The term $\Omega_{\bm{k}}^{s}$ is the Berry curvature [26, 27]
$\bm{\Omega}_{\bm{k}}^{s}$ $=$ $Im[\langle\bm{\nabla}_{k}u_{k}^{s}|\times|\bm{\nabla}_{k}u_{k}^{s}\rangle]$ (4)
$\bm{m}_{\bm{k}}^{s}$ $=$ $-\frac{e}{2\hbar}Im[\langle\bm{\nabla}_{k}u_{k}^{s}|\times(\mathcal{H}_{J}(\bm{k})-\epsilon_{\bm{k}}^{s})|\bm{\nabla}_{k}u_{k}^{s}\rangle]$ (5)
where $|u_{k}^{s}\rangle$ satisfies the equation $\mathcal{H}_{J}(\bm{k})|u_{k}^{s}\rangle=\epsilon_{\bm{k}}^{s}|u_{k}^{s}\rangle$
The general expressions for Berry curvature and orbital magnetic moment for multi-WSMs are [28]
$\bm{\Omega}_{\bm{k}}^{s}=\pm\frac{s}{2}\frac{J\mathit{v}_{F}\alpha_{J}^{2}k_{\perp}^{2J-2}}{\beta_{\bm{k},s}^{3}}\{k_{x},k_{y},Jk_{z}\}$ (6)
$\bm{m}_{\bm{k},s}^{\pm}=\frac{s}{2}\frac{eJ\mathit{v}_{F}\alpha_{J}^{2}k_{\perp}^{2J-2}}{\hbar\beta_{\bm{k},s}^{2}}\{k_{x},k_{y},Jk_{z}\}$ (7)
where $\beta_{\bm{k},s}=\sqrt{\alpha_{J}^{2}k_{\perp}^{2}+\left(k_{z}\mathit{v}_{F}\right)^{2}}$ in the case of mWSMs.
The two equations (2) and (3) can be decoupled to get
$\bm{\dot{r}}=\frac{1}{\hbar D}[\nabla_{\bm{k}}\tilde{\varepsilon}_{\bm{k}}^{s}+e\bm{E}\times\bm{\Omega}_{\bm{k}}^{s})+\frac{e}{\hbar}(\nabla_{\bm{k}}\tilde{\varepsilon}_{\bm{k}}^{s}\cdot\bm{\Omega}_{\bm{k}}^{s})\bm{B}]$ (8)
$\hbar\bm{\dot{k}}=\frac{1}{\hbar D}[-e\bm{E}-\frac{e}{\hbar}\nabla_{\bm{k}}\tilde{\varepsilon}_{\bm{k}}^{s}\times\bm{B}-\frac{e^{2}}{\hbar}(\bm{E}.\bm{B})\bm{\Omega}_{\bm{k}}^{s}]$ (9)
where the factor $D=1+\frac{e}{\hbar}(\bm{\Omega}_{\bm{k}}^{s}\cdot\bm{B})$ modifies the phase space volume [29].
For a given chirality $s=\pm$ of a single Weyl node, the semiclassical Boltzmann equation (SBE) reads as follows
$\frac{\partial\tilde{f}^{s}}{\partial t}+\bm{\dot{k}}.\frac{\partial\tilde{f}^{s}}{\partial\bm{k}}=\frac{\tilde{f}^{s}-\tilde{f}^{s}_{0}}{\tau}$ (10)
[…]
The $\tilde{f}_{0}^{s}(\epsilon_{\bm{k}}^{s})$ can be expanded at low magnetic field as [31]
$\tilde{f}_{0}^{s}(\tilde{\epsilon_{\bm{k}}^{s}})=\tilde{f}_{0}^{s}(\epsilon_{\bm{k}}^{s}-\bm{m}_{\bm{k}}^{s}\cdot\bm{B})$
$\simeq\tilde{f}_{0}^{s}(\epsilon_{\bm{k}}^{s})-\bm{m}_{\bm{k}}^{s}\cdot\bm{B}\frac{\partial\tilde{f}_{0}^{s}(\epsilon_{\bm{k}}^{s})}{\partial\epsilon_{\bm{k}}^{s}}$ (11)
where $\tilde{f}_{0}^{s}(\epsilon_{\bm{k}}^{s})=1/[e^{(\epsilon_{\bm{k}}^{s}-\mu)/k_{B}T}+1]$ with $k_{B}$ the Boltzmann constant, T the temperature, and $\mu$ the chemical potential.
Eq.(10) can be solved by expanding the distribution function as a power series in the electric field as
$\tilde{f}^{s}=\tilde{f}_{0}^{s}+\tilde{f}_{1}^{s}e^{-i\omega t}+\tilde{f}_{2}^{s}e^{-2i\omega t}+.$... (12)
where $\tilde{f}_{1}^{s}$ and $\tilde{f}_{2}^{s}$ are the first- and second-order terms for $\bm{E}$, respectively. The electric current density can be calculated by
$\bm{j}=-\frac{e}{(2\pi)^{3}}\int d^{3}kD\bm{\dot{r}}\tilde{f}^{s}$ (13)
Equation(13) compute the conductivity components under the combined influence of external electric and magnetic fields.
III Linear response of double WSMs
For linear electric field response, we retain only the first two terms of Eq.(12) and substitute Eq.(8) in Eq.(10)
$\frac{1}{\hbar D}[-e\textbf{E}-\frac{e^{2}}{\hbar}(\textbf{E}\cdot\textbf{B})\Omega_{\bm{k}}^{s}]\cdot\frac{\partial\tilde{f}_{0}^{s}}{\partial\bm{k}}-i\omega\tilde{f}_{1}^{s}=-\frac{\tilde{f}_{1}^{s}}{\tau}$ (14)
Solve for $\tilde{f}_{1}^{s}$, we obtain
$\tilde{f}_{1}^{s}=\frac{\tau}{(1-i\omega\tau)}\frac{1}{\hbar D}[e\textbf{E}+\frac{e^{2}}{\hbar}(\textbf{E}\cdot\textbf{B})\Omega_{\bm{k}}^{s}]\cdot\frac{\partial\tilde{f}_{0}^{s}}{\partial\bm{k}}$ (15)
We expand Eq.(15) up to the second order in magnetic field and obtain
$\tilde{f}_{1}^{s}$ $=$ $\frac{\tau}{(1-i\omega\tau)}\biggl{[}e\textbf{E}\cdot\bm{\mathit{v}}_{\bm{k}}^{s}\cdot\frac{\partial\tilde{f}_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}-\frac{e^{2}}{\hbar}(\textbf{B}\cdot\Omega_{\bm{k}}^{s})(\textbf{E}\cdot\bm{\mathit{v}}_{\bm{k}}^{s})\frac{\partial\tilde{f}_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}+\frac{e^{2}}{\hbar}(\textbf{E}\cdot\textbf{B})(\Omega_{\bm{k}}^{s}\cdot\bm{\mathit{v}}_{\bm{k}}^{s})\frac{\partial\tilde{f}_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}-\frac{e}{\hbar}\textbf{E}\cdot\frac{\partial}{\partial\bm{k}}\biggl{(}\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B}\frac{\partial\tilde{f}_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}\biggr{)}$ (16)
- $\frac{e^{3}}{\hbar^{2}}(\textbf{B}\cdot\Omega_{\bm{k}}^{s})(\textbf{E}\cdot\textbf{B})(\Omega_{\bm{k}}^{s}\cdot\bm{\mathit{v}}_{\bm{k}}^{s})\frac{\partial\tilde{f}_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}+\frac{e^{3}}{\hbar^{2}}(\textbf{B}\cdot\Omega_{\bm{k}}^{s})^{2}(\textbf{E}\cdot\bm{\mathit{v}}_{\bm{k}}^{s})\frac{\partial\tilde{f}_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}+\frac{e^{2}}{\hbar^{2}}(\textbf{B}\cdot\Omega_{\bm{k}}^{s})\textbf{E}\cdot\frac{\partial}{\partial\bm{k}}\biggl{(}\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B}\frac{\partial\tilde{f}_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}\biggr{)}$
- $\frac{e^{2}}{\hbar^{2}}\Omega_{\bm{k}}^{s}\cdot\frac{\partial}{\partial\bm{k}}\biggl{(}\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B}\frac{\partial\tilde{f}_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}\biggr{)}\biggr{]}$
From Eqs.(8) and Eq.(16), the expression for current density at time t is given by
$\bm{j}_{1}$ $=$ $-\frac{e}{(2\pi)^{3}}\int d^{3}k\Bigl{[}\bm{\mathit{\tilde{v}}}_{\bm{k}}^{s}+\frac{e}{\hbar}(\Omega_{\bm{k}}^{s}\cdot\bm{\mathit{\tilde{v}}}_{\bm{k}}^{s})\textbf{B}\Bigr{]}\tilde{f}_{1}^{s}$ (17)
- $\frac{e^{2}}{2\pi)^{3}\hbar}\int d^{3}k\textbf{E}\times\Omega_{\bm{k}}^{s}\tilde{f}_{0}^{s}$
The above equation can be expressed in frequency space $\omega$ as
$j_{a}(\omega)=\sigma_{ab}(\omega)E_{b}(\omega)$ (18)
where $\sigma_{ab}(\omega)$ is the frequency dependent conductivity. It is known that the single contribution from the group velocity $\bm{\mathit{\tilde{v}}}_{\bm{k}}^{s}$ or the Berry curvature $\bm{\Omega}_{\bm{k}}^{s}$ form the conventional longitudinal or Hall conductivities. In the presence of the magnetic field, the conductivity $\sigma(\omega)$ consists of the coupling terms between the group velocity $\bm{\mathit{\tilde{v}}}_{\bm{k}}^{s}$ and the Berry curvature $\bm{\Omega}_{\bm{k}}^{s}$ besides the conventional ingredients [see Eq. (17)]. Their combined contributions are triggered by the external magnetic field, and play a crucial role in the electron transport.
III.1 Calculations of longitudinal conductivities components without magnetic field
Substituting Eq.(16) without B terms into the first term of Eq.(17), we get longitudinal conductivities components
$\sigma_{ab}^{(0)}(\omega)=\frac{\tau}{(1-i\omega\tau)}\frac{e^{2}}{(2\pi)^{3}}\int d^{3}k\mathit{v}_{a}^{s}\mathit{v}_{b}^{s}\Bigl{(}-\frac{\partial f_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}\Bigr{)}$ (19)
Refer to caption  Refer to caption  Refer to caption
Figure 1: The dependence of the optical conductivity on the tilt $t_{+}$ at zero B field for (a) single WSM (b)double WSM and (c) triple WSM. The other parameters are taken as $v_{F}=4.13\times 10^{5}$ m / s, $\alpha_{2}=0.009m^{2}/s$, $\alpha_{3}=4.5\times 10^{-11}m^{3}/s$, $\mu=1meV$, and $\tau=10^{-13}s$.
At T=0 K, $\frac{\partial f_{0}^{s}}{\partial\varepsilon_{\bm{k}}^{s}}=\delta(\varepsilon_{\bm{k}}^{s}-\mu)$, we get
for J=1,
$\sigma_{zz}^{0}(\omega)$ $=$ $\frac{\mu^{2}}{\pi\hbar^{2}\mathit{v_{F}}}\frac{\sigma_{D}}{t_{s}^{3}}\Bigl{[}-t_{s}-\frac{1}{2}\ln{\frac{1-t_{s}}{1+t_{s}}}\Bigr{]}$ (20)
$\sigma_{xx}^{0}(\omega)$ $=$ $\sigma_{yy}^{0}(\omega)=\frac{\mu^{2}}{\pi\hbar^{2}\mathit{v_{F}}}\frac{\sigma_{D}}{t_{s}^{3}}\Bigl{[}\frac{t_{s}}{(1-t_{s}^{2})}$ (21)
$+$ $\frac{1}{2}\ln{\frac{1-t_{s}}{1+t_{s}}}\Bigr{]}\frac{1}{2}$
for J=2,
$\sigma_{zz}^{0}(\omega)$ $=$ $\frac{\mu\mathit{v_{F}}}{4\hbar\alpha_{2}}\frac{\sigma_{D}}{t_{s}^{2}}\Bigl{[}1-\sqrt{1-t_{s}^{2}}\Bigr{]}$ (22)
$\sigma_{xx}^{0}(\omega)$ $=$ $\sigma_{yy}^{0}(\omega)=2\frac{\mu^{2}\sigma_{D}}{\pi\mathit{v_{F}}\hbar^{2}{t_{s}^{3}}}\Bigl{[}\frac{t_{s}}{1-t_{s}^{2}}$ (23)
$+$ $\frac{1}{2}\ln\frac{1-t_{s}}{1+t_{s}}\Bigr{]}\frac{1}{2}$
for J=3
$\sigma_{zz}^{o}(\omega)$ $=$ $\frac{\mu^{\frac{2}{3}}\mathit{v_{F}}\sigma_{D}}{\pi^{\frac{1}{2}}\hbar^{\frac{2}{3}}\alpha_{3}^{2/3}}\Bigl{[}\frac{\Gamma(7/6)((-1+t_{s})(1-t_{s}^{2})^{1/3}+(1+t_{s})^{2/3}{}_{2}F_{1}(\frac{2}{3},\frac{4}{3},\frac{5}{3},\frac{2t_{s}}{-1+t_{s}})))}{3^{1/2}t_{s}(1-t_{s}^{2})^{2/3}\Gamma(2/3)}\Bigr{]}$ (24)
$\sigma_{xx}^{0}(\omega)$ $=$ $\sigma_{yy}^{0}(\omega)=3\frac{\mu^{2}}{\pi\hbar^{2}\mathit{v_{F}}}\frac{\sigma_{D}}{t_{s}^{3}}\Bigl{[}\frac{t_{s}}{(1-t_{s}^{2})}+\frac{1}{2}\ln\frac{1-t_{s}}{1+t_{s}}\Bigr{]}\frac{1}{2}$ (25)
Refer to caption  Refer to caption  Refer to caption
Figure 2: The frequency dependence of optical conductivity at t=0.5 at zero B-field for (a) single WSM (b)double WSM and (c) triple WSM. The other parameters are the same as those of Fig.(1)
where $\sigma_{D}=\frac{e^{2}\tau}{(1-i\omega\tau)2\pi\hbar}$ is Drude frequency complex conductivity. The $\sigma_{zz}^{0}(\omega)$ are modified compared to single WSM while $\sigma_{xx}^{0}(\omega)$ have the same form as in the case of single WSMs. Eqs.(20) and (21). In the limit $t\rightarrow$0, $\sigma_{xx}^{0}(\omega)=\sigma_{yy}^{0}(\omega)=J\frac{\mu^{2}}{3\pi\hbar^{2}v_{F}}\sigma_{D}$,J=1 $\sigma_{zz}^{0}(\omega)=\frac{\mu^{2}}{3\pi\hbar^{2}v_{F}}\sigma_{D}$, J=2 $\sigma_{zz}^{0}(\omega)=\frac{\mu v_{F}}{8\hbar}\sigma_{D}$ and J=3 $\sigma_{zz}^{0}(\omega)=\sqrt{\frac{3}{\pi}}\frac{\Gamma[7/6]}{\Gamma[2/3]}\frac{\mu^{2/3}v_{F}}{5\hbar^{2/3}\alpha_{3}^{2/3}}\sigma_{D}$. Note that $\sigma_{xx}^{0}(\omega)\propto J\mu^{2}$ while $\sigma_{xx}^{0}(\omega)\propto\frac{\mu^{2/J}}{\alpha_{J}^{2/J}\hbar^{2/J}}$ obeys the power-laws in chemical potential. We find from Eqs.(20) to (25), the conductivity components are even with respect to the tilt parameter $t_{s}$. Thus, the total conductivity elements will be twice the contribution from single Weyl nodes. Fig.(1) shows a comparison between different m-WSMs on the tilt parameter $t_{+}$ dependence of $\sigma^{(0)}_{xx}(\omega)$ and $\sigma^{(0)}_{zz}(\omega)$. In addition, the frequency dependence of the conductivity $\sigma_{aa}^{(0)}(\omega)(a=x,z)$ exhibit Drude-type behavior as shown in Fig.(2).
Refer to caption  Refer to caption  Refer to caption
Figure 3: The dependence of the optical conductivity on the tilt $t_{+}=0.5$. for (a) single WSM (b)double WSM and (c) triple WSM. The other parameters are the same as those of Fig.(1)
III.2 Calculations of conductivities components linear in magnetic field
Substituting Eq.(16) with B terms up to first order into the first term of Eq.(17), we get conductivities components
$\sigma_{ab}^{(B)}(\omega)=\sigma_{ab}^{(B,\Omega)}(\omega)+\sigma_{ab}^{(B,m)}(\omega)$ (26)
where
$\sigma_{ab}^{(B,\Omega)}(\omega)$ $=$ $\frac{\tau}{\hbar(1-i\omega\tau)}\frac{e^{3}}{(2\pi)^{3}}\int d^{3}k[(\mathit{v}_{a}^{s}B_{b}+\mathit{v}_{b}^{s}B_{a})(\Omega_{\bm{k}}^{s}\cdot\bm{\mathit{v}}_{\bm{k}}^{s})-\mathit{v}_{a}^{s}\mathit{v}_{b}^{s}(\Omega_{\bm{k}}^{s}\cdot\textbf{B})]\Bigl{(}-\frac{\partial f_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}\Bigr{)}$ (27)
$\sigma_{ab}^{(B,m)}(\omega)$ $=$ $\frac{\tau}{\hbar(1-i\omega\tau)}\frac{e^{2}}{(2\pi)^{3}}\int d^{3}k\Bigl{[}\frac{\partial\mathit{v}_{a}^{s}}{\partial k_{b}}(\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B})-\frac{\partial(\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B})}{\partial k_{a}}\mathit{v}_{b}^{s}\Bigr{]}\Bigl{(}-\frac{\partial f_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}\Bigr{)}$ (28)
We can easily check from Eqs.(27) and (28) that $\sigma_{ab}^{(B)}(\omega)=\sigma_{ba}^{(B)}(\omega)$. The system possesses the time-reversal symmetry without tilt term and therefore conductivities will vanish [22, 32]. However, the time-reversal symmetry is broken for a finite value of tilt $t_{s}\neq 0$. The first term and the second term in Eq. (27) is related to the Berry curvature $\Omega$ and the orbital magnetic moment respectively and has been studued in details for isotropic WSMs [22]. We will explore our study for double and triple-WSMs case. In the following, the detailed analysis of Eq. (27) is given by considering the magnetic field B perpendicular and parallel to the tilt direction $t_{s}$.
Refer to caption  Refer to caption  Refer to caption
Figure 4: The frequency dependence of optical conductivity at t=0.5 and B = 1 T for (a) single WSM (b)double WSM and (c) triple WSM. The other parameters are the same as those of Fig.(1)
Case-I When $\textbf{B}\|t_{s}\|z$
In this case, Hall conductivities components are zero, and one can get the following expressions for longitudinal components.
for J=1
$\sigma_{zz}^{(B,\Omega)}$ $=$ $\frac{\mathit{v_{F}}\sigma_{1}^{(B)}s}{4\pi^{2}\hbar}\Bigl{[}\frac{2(3-5t_{s}^{2}-3t_{s}^{4})}{3t_{s}^{2}}-\frac{(t_{s}^{2}-1)^{2}}{t_{s}^{4}}\ln\frac{1-t_{s}}{1+t_{s}}\Bigr{]}$
$\sigma_{xx}^{(B,\Omega)}$ $=$ $\frac{\mathit{v_{F}}\sigma_{1}^{(B)}s}{4\pi^{2}\hbar}\Bigl{[}\frac{2t_{s}^{2}-3}{3t_{s}^{3}}-\frac{1-t_{s}^{2}}{2t_{s}^{4}}\ln\frac{1-t_{s}}{1+t_{s}}B\Bigr{]}$ (30)
for J=2
$\sigma_{zz}^{(B,\Omega)}$ $=$ $\frac{2\mathit{v_{F}}\sigma_{1}^{(B)}s}{4\pi^{2}\hbar}\Bigl{[}\frac{2(3-5t_{s}^{2}-3t_{s}^{4})}{3t_{s}^{2}}-\frac{(t_{s}^{2}-1)^{2}}{t_{s}^{4}}\ln\frac{1-t_{s}}{1+t_{s}}\Bigr{]}$
$\sigma_{xx}^{(B,\Omega)}=\frac{\mu\alpha_{2}\sigma_{1}^{(B)}s}{\pi\mathit{v_{F}\hbar^{2}}}\Bigl{[}\frac{(-4+4\sqrt{1-t_{s}^{2}}+t_{s}^{2}(3-\sqrt{1-t_{s}^{2}})}{t_{s}^{5}}\Bigr{]}$
for J= 3
$\sigma_{zz}^{(B,\Omega)}$ $=$ $\frac{3\mathit{v}_{F}\sigma_{1}^{(B)}s}{4\pi^{2}\hbar}\Bigl{[}\frac{2(-3+5t_{s}^{2}+3t_{s}^{4})}{3t_{s}^{3}}-\frac{(-1+t_{s}^{2})^{2}}{t_{s}^{4}}\ln\frac{1-t_{s}}{1+t_{s}}\Bigr{]}$ (33)
$\sigma_{xx}^{(B,\Omega)}$ $=$ $\frac{\mu^{\frac{4}{3}}\alpha_{3}^{\frac{2}{3}}\sigma_{1}^{(B)}s}{4\pi^{\frac{3}{2}}\mathit{v}_{F}\hbar^{\frac{7}{3}}}\Bigl{[}-\frac{(1152(3)^{\frac{1}{2}}\Gamma(\frac{17}{6})({{}_{2}F_{1}}(\frac{7}{3},\frac{8}{3},\frac{16}{3},\frac{2t_{s}}{-1+t_{s}})-{{}_{2}F_{1}}(\frac{7}{3},\frac{11}{3},\frac{19}{3},\frac{2t_{s}}{-1+t}))}{143(1-t_{s})^{\frac{7}{3}}\Gamma(\frac{10}{3})}\Bigr{]}$ (34)
where $\sigma_{1}^{(B)}=\frac{e^{3}\tau B}{(1-i\omega\tau)2\hbar}$.
So, we can rewrite the above components follow the power laws $\sigma_{zz}^{(B,\Omega)}(\omega)\propto J$ and $\sigma_{xx}^{(B,\Omega)}(\omega)\propto\frac{\mu^{2\bigl{(}1-\frac{1}{J}\bigr{)}}\alpha_{J}^{2/J}}{\hbar^{\bigl{(}3-\frac{2}{J}\bigr{)}}}$.
Via the similar calculation, Eq.(28) becomes for J=1
$\sigma_{zz}^{(B,m)}(\omega)$ $=$ $\frac{\mathit{v_{F}}\sigma_{1}^{(B)}s}{4\pi^{2}\hbar}\Bigl{[}\frac{2(-3+5t_{s}^{2})}{3t_{s}^{3}}-\frac{(t_{s}^{2}-1)^{2}}{t_{s}^{4}}\ln\frac{1-t_{s}}{1+t_{s}}\Bigr{]}$
$\sigma_{xx}^{(B,m)}(\omega)$ $=$ $\frac{\mathit{v_{F}}\sigma_{1}^{(B)}s}{4\pi^{2}\hbar}\Bigl{[}\frac{(3-8t_{s}^{2})}{3t_{s}^{3}}-\frac{1-3t_{s}}{2t_{s}^{4}}\ln\frac{1-t_{s}}{1+t_{s}}\Bigr{]}$ (36)
for J=2
$\sigma_{zz}^{(B,m)}(\omega)$ $=$ $\frac{\sigma_{1}^{(B)}s}{3\pi^{2}}\Bigl{[}\frac{-3+5t_{s}^{2}}{3t_{s}^{3}}-\frac{(1-t_{s}^{2})^{2}}{2t_{s}^{4}}\ln\frac{1-t_{s}}{1+t_{s}}\Bigr{]}$ (37)
$\sigma_{xx}^{(B,m)}(\omega)$ $=$ $\frac{\mu\alpha_{2}\sigma_{1}^{(B)}s}{\pi\mathit{v}_{F}\hbar^{2}}\Bigl{[}\frac{(-4-2t_{s}^{4}+4\sqrt{1-t_{s}^{2}}+t_{s}^{2}(7-5\sqrt{1-t_{s}^{2}})}{t_{s}^{5}\sqrt{1-t_{s}^{2}}}\Bigr{]}$
for J=3
$\sigma_{zz}^{(B,m)}(\omega)$ $=$ $\frac{\mathit{v_{F}}\sigma_{1}^{(B)}s}{2\pi^{2}\hbar}\Bigl{[}\frac{(-3+5t_{s}^{2})}{t_{s}^{3}}-\frac{3(-1+t_{s}^{2})^{2}}{2}\ln{\frac{1-t_{s}}{1+t_{s}}}\Bigr{]}$ (39)
$\sigma_{xx}^{(B,m)}(\omega)$ $=$ $\frac{7\alpha_{3}^{\frac{2}{3}}\mu^{\frac{4}{3}}\sigma_{1}^{(B)}s}{\pi^{\frac{1}{2}}\mathit{v}_{F}\hbar^{\frac{4}{3}}}\Bigl{[}\frac{\frac{5(-65-39t_{s}+93t_{s}^{2})(1-t_{s})^{\frac{8}{3}}}{t_{s}^{3}(1+t_{s})^{\frac{2}{3}}}+243\frac{(1-t_{s})^{5/3}}{(1+t_{s})^{2/3}}-\frac{5(1+t_{s})(-65+126t_{s}^{2}){}_{2}F_{1}(\frac{7}{3},\frac{8}{3},\frac{10}{3},\frac{2t_{s}}{-1+t_{s}})}{t_{s}^{3}}}{243(3)^{\frac{1}{2}}(1-t_{s})^{\frac{7}{3}}(1+t_{s})\Gamma(\frac{13}{6})(\frac{7}{3})}\Bigr{]}$
where $\sigma_{1}^{(B)}=\frac{e^{3}\tau B}{(1-i\omega\tau)2\hbar}$.
So, we can summarize the above components follow the power laws $\sigma_{zz}^{(B,m)}(\omega)\propto J$ and $\sigma_{xx}^{(B,m)}(\omega)\propto\frac{\mu^{2\bigl{(}1-\frac{1}{J}\bigr{)}}\alpha_{J}^{2/J}}{\hbar^{2(\bigl{(}1-\frac{1}{J}\bigr{)}}}$.
Case-II For $\textbf{B}\perp t_{s}\perp z$
We can represent the magnetic field in the x-y plane as $\textbf{B}=B(\hat{i}\cos\gamma+\hat{j}\sin\gamma)$ where $\gamma$ is the angle between magnetic field B and x-axis. In this case, longitudinal components of conductivities are zero and we get the following expressions for planer Hall conductivities [22, 33]
$\sigma_{xz}^{(B)}(\omega)=[\sigma^{(B,\Omega)}(\omega)+\sigma^{(B,m)}(\omega)]\cos\gamma$ (41)
$\sigma_{yz}^{(B)}(\omega)=[\sigma^{(B,\Omega)}(\omega)+\sigma^{(B,m)}(\omega)]\sin\gamma$ (42)
$\sigma^{(B,\Omega)}_{xz}(\omega)=\frac{\tau}{\hbar(1-i\omega\tau)}\frac{e^{3}}{(2\pi)^{3}}\int d^{3}k[\mathit{v}_{z}^{s}B\cos\gamma(\mathit{v}_{x}^{s}\Omega_{ky}+\mathit{v}_{z}^{s}\Omega_{kz})-\mathit{v}_{z}^{s}B\sin\gamma\mathit{v}_{x}^{s}\Omega_{ky}]\Bigl{(}-\frac{\partial f_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}\Bigr{)}$ (43)
$\sigma^{(B,m)}_{xz}(\omega)=\frac{\tau}{\hbar(1-i\omega\tau)}\frac{e^{2}}{(2\pi)^{3}}\int d^{3}k\Bigl{[}B\cos\gamma\Bigl{(}\frac{\partial\mathit{v}_{x}^{s}}{\partial k_{z}}m_{kx}^{s}-v_{z}^{s}\frac{\partial\mathit{m}_{kx}^{s}}{\partial k_{z}}\Bigr{)}-B\sin\gamma\Bigl{(}\frac{\partial\mathit{v}_{x}^{s}}{\partial k_{z}}m_{ky}^{s}-v_{z}^{s}\frac{\partial\mathit{m}_{ky}^{s}}{\partial k_{x}}\Bigr{)}\Bigr{]}\Bigl{(}-\frac{\partial f_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}\Bigr{)}$ (44)
The second term of both the above equations contributes to zero.
for J=1
$\sigma^{(B,\Omega)}$ $=$ $\frac{\mathit{v_{F}}\sigma_{1}^{(B)}s}{4\pi^{2}\hbar}\Bigl{[}\frac{-3+5t_{s}^{2}-6t_{s}^{4}}{3t_{s}^{3}}-\frac{(1-t_{s}^{2})^{2}}{2t_{s}^{4}}\ln\frac{1-t_{s}}{1+t_{s}}\Bigr{]}$
$\sigma^{(B,m)}=\frac{\mathit{v_{F}}\sigma_{1}^{(B)}s}{4\pi^{2}\hbar}\Bigl{[}-\frac{2t_{s}^{2}-3}{3t_{s}^{3}}+\frac{1-t_{s}^{2}}{2t_{s}^{4}}\ln\frac{1-t_{s}}{1+t_{s}}\Bigr{]}$ (46)
for J=2
$\sigma^{(B,\Omega)}$ $=$ $2\frac{\mathit{v_{F}}\sigma_{1}^{(B)}s}{4\pi^{2}\hbar}\Bigl{[}\frac{-3t_{s}+5t_{s}^{3}-6t_{s}^{5}}{3t_{s}^{4}}-\frac{(1-t_{s}^{2})^{2}}{2t_{s}^{4}}\ln\frac{1-t_{s}}{1+t_{s}}\Bigr{]}$
$\sigma^{(B,m)}=3\frac{\mathit{v_{F}}\sigma_{1}^{(B)}s}{6\pi^{2}\hbar}\Bigl{[}\frac{(3-2t_{s}^{2})}{3t_{s}^{3}}+\frac{(1-t_{s}^{2})}{2t_{s}^{4}}\ln\frac{1-t_{s}}{1+t_{s}}\Bigr{]}$ (48)
for J=3
$\begin{split}\sigma^{(B,\Omega)}=&3\frac{\mathit{v_{F}}\sigma_{1}^{(B)}s}{4\pi^{2}\hbar}\Bigl{[}\frac{(-3t_{s}+5t_{s}^{3}-6t_{s}^{5})}{3t_{s}^{4}}\\ &-\frac{(-1+t_{s}^{2})^{2}}{2t_{s}^{4}}\ln{\frac{1-t_{s}}{1+t_{s}}}\Bigr{]}\end{split}$ (49)
$\sigma^{(B,m)}=3\frac{\mathit{v_{F}}\sigma_{1}^{(B)}s}{4\pi^{2}\hbar}\Bigl{[}\frac{(3-2t_{s}^{2})}{3t_{s}^{3}}-\frac{(-1+t_{s}^{2})}{2t_{s}^{4}}\ln{\frac{1-t_{s}}{1+t_{s}}}\Bigr{]}$ (50)
where $\sigma_{0}^{(B)}=\frac{e^{3}\tau B}{(1-i\omega\tau)2\hbar}$. So, we can rewrite the above components follow the power laws $\sigma^{(B,\Omega)}\propto\frac{J}{\hbar}$ and $\sigma^{(B,m)}\propto\frac{J}{\hbar}$.
From Eqs. (33)–(50), we notice that the B-linear magnetoconductivity in tilted mult-Weyl semimetals is independent of Fermi energy $\mu$, and the odd function of $t_{s}$ as in the case of single WSMs [22]. According to the Nielsen-Ninomiya theorem [5, 5], the Weyl nodes with opposite chirality always appears in pairs. The total magnetoconductivity of the system is the sum of all the Weyl nodes. Therefore, for the case of $t_{+}=t_{-}$, where the tilt inversion symmetry is broken, the contribution of the Weyl node to the magnetoconductivity has the opposite sign for the opposite (B) chirality, giving rise to $\sigma_{ab}(\omega)$ = 0. Whereas, for the case of $t_{+}=-t_{-}$, where the tilt inversion symmetry is unbroken, each Weyl node produces an identical contribution to the magnetoconductivity, and so the nonzero magnetoconductivity emerges for this case. Figure (3) shows the B-linear magnetoconductivity as a function of the tilt $t_{+}$ . It is observed that the contribution from the orbital magnetic moment is always negative and decreases with increasing $t_{+}$, which partially cancels the contribution of the Berry curvature to magnetoconductivity [22]. The total magnetoconductivity tends to slow down at the large $t_{+}$ . Figure (4) shows the B-linear magnetoconductivity as a function of the THz incident light.
III.3 Calculations of quadratic-B contribution to the conductivity $\sigma_{ab}^{(B^{2})}$
Substituting Eq.(16) with B terms up to second order into the first term of Eq.(17), we get conductivities components
$\sigma_{ab}^{(B^{2})}(\omega)=\sigma_{ab}^{(B^{2},\Omega)}(\omega)+\sigma_{ab}^{(B^{2},m)}(\omega)$ (51)
where
$\begin{split}\sigma_{ab}^{(B^{2},\Omega)}(\omega)=&\frac{\tau}{\hbar^{2}(1-i\omega\tau)}\frac{e^{4}}{(2\pi)^{3}}\int d^{3}k[\mathit{v}_{a}^{s}\mathit{v}_{b}^{s}(\Omega_{\bm{k}}^{s}\cdot\textbf{B})^{2}-(\mathit{v}_{a}^{s}B_{b}+\mathit{v}_{b}^{s}B_{a})(\Omega_{\bm{k}}^{s}\cdot\bm{\mathit{v}}_{\bm{k}}^{s})(\Omega_{\bm{k}}^{s}\cdot\textbf{B})\\ &+B_{a}B_{b}(\Omega_{\bm{k}}^{s}\cdot\bm{\mathit{v}}_{\bm{k}}^{s})^{2}]\Bigl{(}-\frac{\partial f_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}\Bigr{)}\end{split}$ (52)
$\begin{split}\sigma_{ab}^{(B^{2},m)}(\omega)=&\frac{\tau}{\hbar^{2}(1-i\omega\tau)}\frac{e^{3}}{(2\pi)^{3}}\int d^{3}k\Bigl{[}\frac{\partial}{\partial\textbf{k}}\cdot[\mathit{v}_{a}^{s}B_{b}\Omega_{\bm{k}}^{s}](\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B})-\frac{\partial[\mathit{v}_{a}^{s}(\Omega_{\bm{k}}^{s}\cdot\textbf{B})]}{\partial k_{b}}(\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B})+\frac{\partial[B_{a}^{s}(\Omega_{\bm{k}}^{s}\cdot\mathit{\bm{v}}_{\bm{k}}^{s})]}{\partial k_{b}}(\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B})+\\ &\frac{\partial(\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B})}{\partial k_{a}}(\Omega_{\bm{k}}^{s}\cdot\textbf{B})\mathit{v}_{b}^{s}-\frac{\partial(\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B})}{\partial k_{a}}(\Omega_{\bm{k}}^{s}\cdot\mathit{\bm{v}}_{\bm{k}}^{s})B_{b}^{s}-\frac{\partial^{2}(\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B})}{e\partial k_{a}\partial k_{b}}(\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B})-B_{a}\frac{\partial(\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B})}{\partial\bm{k}}\cdot\Omega_{\bm{k}}^{s}\mathit{v}_{b}^{s}\Bigr{]}\Bigl{(}-\frac{\partial f_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}\Bigr{)}\end{split}$ (53)
$\sigma_{ab}^{(B^{2},\Omega)}(\omega)$ and $\sigma_{ab}^{(B^{2},m)}(\omega)$ Berry-curvature $\Omega_{\bm{k}}^{s}$ and the orbital magnetic moment $\textbf{m}_{\bm{k}}^{s}$ dependent conductivities respectively. Consider the following two cases
Case-I When $\textbf{B}\|t_{s}\|z$
In this case, Hall conductivities components are zero and one can get the following expressions for longitudinal components
for J=1
$\sigma_{zz}^{(B^{2},\Omega)}(\omega)$ $=$ $\frac{\mathit{v}_{F}^{3}}{15\pi\mu^{2}}\sigma_{1}^{(B^{2})}$ (54)
$\sigma_{xx}^{(B^{2},\Omega)}(\omega)$ $=$ $\frac{\mathit{v}_{F}^{3}}{120\pi\mu^{2}}\sigma_{1}^{(B^{2})}$ (55)
for J=2
$\sigma_{zz}^{(B^{2},\Omega)}(\omega)$ $=$ $\frac{5\mathit{v}_{F}\alpha_{2}^{2}}{32\mu\hbar}\sigma_{1}^{(B^{2})}$ (56)
$\sigma_{xx}^{(B^{2},\Omega)}(\omega)$ $=$ $\frac{\alpha_{2}^{2}\sigma_{1}^{(B^{2})}}{15\pi\mathit{v}_{F}\hbar^{2}}\Bigl{[}\frac{-2(15-25t_{s}^{2}+8t_{s}^{4})}{t_{s}^{6}}$ (57)
- $15\frac{(1-t_{s}^{2})^{2}}{t_{s}^{7}}\ln\frac{1-t_{s}}{1+t_{s}}\Bigr{]}$
for J=3
$\sigma_{zz}^{(B^{2},\Omega)}(\omega)=\frac{\mathit{v}_{F}\alpha_{3}^{\frac{2}{3}}\sigma_{1}^{(B^{2})}}{8\pi\mu^{\frac{2}{3}}\hbar^{\frac{4}{3}}}\Bigl{[}\frac{59049\Gamma{(\frac{11}{33})^{2}}{}_{2}F_{1}{(\frac{1}{3},\frac{11}{3},\frac{22}{3},\frac{2t_{s}}{-1+t_{s}})}}{34580{(\frac{1}{2}-\frac{t_{s}}{2})^{\frac{1}{3}}}\Gamma{(\frac{1}{3})}}\Bigr{]}$ (58)
$\begin{split}\sigma_{xx}^{(B^{2},\Omega)}(\omega)=&\frac{\mu^{\frac{2}{3}}\alpha^{\frac{4}{3}}\pi^{\frac{1}{2}}\sigma_{1}^{(B^{2})}}{8\mathit{v}_{F}\hbar^{\frac{8}{3}}}\frac{1}{{27{(3)^{\frac{1}{2}}}{(1-t_{s})^{\frac{2}{3}}}t_{s}^{4}{(-1+t_{s}^{2})}\Gamma{(\frac{5}{6})}\Gamma{(\frac{17}{3})}}}\bigg{[}6272(1+t_{s})\bigg{\{}\big{(}1190+t_{s}(1530+t_{s}(15-t_{s}(335+24t_{s}))\big{)}\\ &{}_{2}F_{1}{\Big{(}\frac{-1}{3},\frac{10}{3},\frac{14}{3},\frac{2t_{s}}{-1+t_{s}}\Big{)}}+5(1+t_{s})(-238+t_{s}(-34+t_{s}(53+3t_{s})))_{2}F_{1}{\Big{(}\frac{-1}{3},\frac{13}{3},\frac{14}{3},\frac{2t_{s}}{-1+t_{s}}\Big{)}}\bigg{\}}\bigg{]}\end{split}$ (59)
So, we can rewrite the above components follow the power laws $\sigma_{zz}^{(B^{2},\Omega)}(\omega)\propto\frac{\alpha_{J}^{\frac{2}{J}}}{\mu^{\frac{2}{J}}\hbar^{(2-\frac{2}{J})}}$.
Considering the effect of the orbital magnetic moment, we have
for J=1
$\sigma_{zz}^{(B^{2},m)}(\omega)$ $=$ $\frac{\mathit{v}_{F}^{3}}{120\pi\mu^{2}}\sigma_{1}^{(B^{2})}\Bigl{[}-3+5t_{s}^{2}\Bigr{]}$ (60)
$\sigma_{xx}^{(B^{2},m)}(\omega)$ $=$ $-\frac{\mathit{v}_{F}^{3}}{120\pi\mu^{2}}\sigma_{1}^{(B^{2})}$ (61)
for J=2
$\sigma_{zz}^{(B^{2},m)}(\omega)$ $=$ $\frac{\mathit{v}_{F}\alpha_{2}}{8\mu\hbar}\sigma_{1}^{(B^{2})}$ (62)
$\sigma_{xx}^{(B^{2},m)}(\omega)$ $=$ $\frac{2\alpha_{2}^{2}}{5\pi\hbar^{2}\mathit{v}_{F}}\sigma_{1}^{(B^{2})}\Bigl{[}\frac{(30-20t_{s}^{2}+t_{s}^{4})}{t_{s}^{6}}$ (63)
$+$ $\frac{5}{2}\frac{(6-6t_{s}^{2}+t_{s}^{4})}{t_{s}^{7}}\ln\frac{1-t_{s}}{1+t_{s}}]$
for J=3
$\sigma_{zz}^{(B^{2},m)}(\omega)=\frac{\mathit{v}_{F}\alpha_{3}^{\frac{2}{3}}\sigma_{1}^{(B^{2})}}{8\pi\mu^{\frac{2}{3}}\hbar^{\frac{4}{3}}}\frac{2^{\frac{1}{3}}(-(-1+t_{s})^{\frac{1}{3}}\Gamma{(\frac{-1}{3})^{2}}}{{729t_{s}^{5}(-(-1+t_{s}))^{\frac{2}{3}}\Gamma{(\frac{7}{3})}}}\bigg{[}-\big{\{}728+3t_{s}\big{(}104+3t_{s}(-117+t_{s}(-41+3t_{s}(1+t_{s})(15+2t_{s})))\big{)}\big{\}}\\ {}_{2}F_{1}{\Big{(}\frac{1}{3},\frac{5}{3},\frac{7}{3},\frac{2t_{s}}{-1+t_{s}}\Big{)}}-{(1+t_{s})\big{\{}-728+3t_{s}\big{(}-208+t_{s}(143+3t_{s}(34+t_{s}(-11+6t_{s})))\big{)}\big{\}}}{}_{2}F_{1}{\Big{(}\frac{4}{3},\frac{5}{3},\frac{7}{3},\frac{2t_{s}}{-1+t_{s}}\Big{)}}\Bigr{]}$ (64)
$\sigma_{xx}^{(B^{2},m)}(\omega)=\frac{\mu^{\frac{2}{3}}\alpha_{3}^{\frac{4}{3}}\sigma_{1}^{(B^{2})}}{8\pi^{\frac{1}{2}}\hbar^{\frac{8}{3}}\mathit{v}_{F}}\frac{-16\Gamma{(\frac{7}{6})}}{5(3)^{\frac{1}{2}}(1-t_{s})^{\frac{2}{3}}t_{s}^{6}(-1+t_{s}^{2})\Gamma{(\frac{5}{3})}}\bigg{[}\big{\{}5236+t_{s}\big{(}1309+t_{s}(-6762+t_{s}(-1393+2t_{s}(982+\\ 3(41-10t_{s})t_{s})))\big{)}\big{\}}{}_{2}F_{1}{\Big{(}\frac{-1}{3},\frac{4}{3},\frac{8}{3},\frac{2t_{s}}{-1+t_{s}}\Big{)}}+4(-1309+1631t_{s}^{2}-442t_{s}^{4}+12t_{s}^{6}){{}_{2}F_{1}{\Big{(}\frac{-1}{3},\frac{7}{3},\frac{8}{3},\frac{2t_{s}}{-1+t_{s}}\Big{)}}}\Big{]}$ (65)
where $\sigma_{1}^{(B^{2})}=\frac{e^{4}\tau B^{2}}{\pi\hbar(1-i\omega t)}$.
So, we can rewrite the above components follow the power laws $\sigma_{zz}^{(B^{2},m)}(\omega)\propto\frac{\alpha_{J}^{\frac{2}{J}}}{\mu^{\frac{2}{J}}}$ and $\sigma_{xx}^{(B^{2},m)}(\omega)\propto\frac{\alpha_{J}^{\frac{4}{J}}}{\mu^{\frac{2}{J}}\hbar^{(2-\frac{2}{J})}}$.
Obviously, from Eqs.(55) and (61), $\sigma_{xx}^{(B^{2},\Omega)}(\omega)$ and $\sigma_{xx}^{(B^{2},m)}(\omega)$, thus the total conductivity $\sigma_{xx}^{(B^{2})}(\omega)$ is equal to zero in the case of single-WSM [22]. For double WSM the total conductivity is enhanced(suppressed) below(above) the critical value $t_{0}$ of the tilt parameter as shown in Fig.(5a). In case of triple WSMs, the total conductivity is enhanced due to the orbital contribution(see Fig.(5b). Further, in the case of single WSM, the total conductivity component $\sigma_{zz}^{(B^{2})}(\omega)$ is suppressed(enhanced) below(above) the value of $t_{0}$ as shown in Fig.(6a)(see ref.[22]). For double WSM, the conductivity component $\sigma_{zz}^{(B^{2})}(\omega)$ the total conductivity is enhanced with tilt parameter $t_{0}$ as shown in Fig.(6b). For triple-WSM, the conductivity component $\sigma_{zz}^{(B^{2})}(\omega)$ the total conductivity is suppressed with tilt parameter $t_{0}$ as shown in Fig.(6c).
Refer to caption  Refer to caption
Figure 5: The dependence of the optical conductivity on the tilt $t_{+}=0.5$ for the case of B $\parallel t_{s}$. for (a) double WSM and (b) triple WSM. The other parameters are the same as those of Fig.(1)
Case-II In this case both longitudinal components as well as planer Hall conductivities are non-zero. The longitudinal components have the following expressions
for J=1
$\sigma_{xx}^{(B^{2},\Omega)}(\omega)$ $=$ $\frac{\mathit{v}_{F}^{3}\sigma_{1}^{(B^{2})}}{120\pi\mu^{2}}\Bigl{[}(8+13t_{s}^{2})\cos^{2}\gamma+\sin^{2}\gamma\Bigr{]}$ (66)
$\sigma_{zz}^{(B^{2},\Omega)}(\omega)$ $=$ $\frac{\mathit{v}_{F}^{3}}{120\pi\mu^{2}}\sigma_{1}^{(B^{2})}(1+7t_{s}^{2})$ (67)
for J=2
$\sigma_{xx}^{(B^{2},\Omega)}(\omega)$ $=$ $\frac{\alpha_{2}\mathit{v}_{F}}{256\pi^{2}\mu\hbar^{2}}\sigma_{1}^{(B^{2})}[(31+16t_{s}^{2})\cos^{2}\gamma+5\sin^{2}\gamma]$
$\sigma_{zz}^{(B^{2},\Omega)}(\omega)$ $=$ $\frac{\mathit{v}_{F}^{3}\sigma_{1}^{(B^{2})}}{60\pi\mu^{2}}(1+7t_{s}^{2})$ (69)
for J=3
$\begin{split}\sigma_{xx}^{(B^{2},\Omega)}(\omega)&=\frac{\mathit{v}_{F}\alpha_{3}^{\frac{2}{3}}\sigma_{1}^{(B^{2})}}{8\pi\hbar^{\frac{4}{3}}\mu^{\frac{2}{3}}}\Bigg{[}\cos^{2}\gamma\frac{1}{670208(1-t_{s})^{\frac{1}{3}}t_{s}^{4}}27\pi\bigg{\{}\big{(}4860-10917t_{s}^{2}+74328t_{s}^{4}+30464t_{s}\big{)}{}_{2}F_{1}{\Big{(}\frac{1}{3},\frac{2}{3},3,\frac{2t_{s}}{-1+t_{s}}\Big{)}}+(1+t_{s})\\ &\big{(}-40860+7677t_{s}^{2}+11288t_{s}^{4}\big{)}{{}_{2}F_{1}\big{(}\frac{4}{3},\frac{3}{2},3,\frac{2t_{s}}{-1+t_{s}}\big{)}}\bigg{\}}+\sin^{2}\gamma\frac{59049\Gamma(\frac{11}{3})^{2}{{}_{2}F_{1}\Big{(}\frac{1}{3},\frac{11}{3},\frac{22}{3},\frac{2t_{s}}{-1+t_{s}}\big{)}}}{276640\big{(}\frac{1}{2}-\frac{t_{s}}{2}\Big{)}^{\frac{1}{3}}\Gamma\big{(}\frac{1}{3}\big{)}}\bigg{]}\end{split}$ (70)
$\sigma_{zz}^{(B^{2},\Omega)}(\omega)=\frac{\mathit{v}_{F}^{3}\sigma_{1}^{(B^{2})}}{40\pi\mu^{2}}(1+7t_{s}^{2})$ (71)
So, we can rewrite the above components follow the power laws $\sigma_{zz}^{(B^{2},\Omega)}(\omega)\propto\frac{J}{\mu^{2}}$ and $\sigma_{xx}^{(B^{2},\Omega)}(\omega)\propto\frac{\alpha_{J}^{\frac{2}{J}}}{\mu^{\frac{2}{J}}\hbar^{2(1-\frac{1}{J})}}$.
The off-diagonal components of conductivities
for J=1
$\sigma_{xy}^{(B^{2},\Omega)}(\omega)=\frac{\mathit{v}_{F}^{3}}{120\pi\mu^{2}}\sigma_{1}^{(B^{2})}(7+13t_{s}^{2})\sin\gamma\cos\gamma$ (72)
for J=2
$\sigma_{xy}^{(B^{2},\Omega)}(\omega)=\frac{\mathit{v}_{F}\alpha_{2}}{128\hbar\mu}\sigma_{1}^{(B^{2})}(13+8t_{s}^{2})\sin\gamma\cos\gamma$ (73)
for J=3
$\sigma_{xy}^{(B^{2},\Omega)}(\omega)$ $=$ $\frac{\mathit{v}_{F}\alpha_{3}^{\frac{2}{3}}\sigma_{1}^{(B^{2})}}{8\pi^{\frac{3}{2}}\mu^{\frac{2}{3}}\hbar^{\frac{4}{3}}}\sin\gamma\cos\gamma\Biggl{[}\frac{1}{189(1-t_{s})^{\frac{1}{3}}}{\bigg{\{}15309(-1+t_{s})^{2}\Gamma{\Big{(}\frac{2}{3}\Big{)}}\Gamma{\Big{(}\frac{5}{6}\Big{)}}{{}_{2}F_{1}{\Big{(}\frac{1}{3},\frac{5}{3},\frac{7}{3},\frac{2t_{s}}{-1+t_{s}}\Big{)}}}}$ (74)
$+$ $\frac{1}{t_{s}^{4}\Gamma{(\frac{1}{6})\Gamma{(\frac{1}{3})}}}2{(3)^{\frac{1}{2}}}\pi^{2}\big{\{}\big{(}-130-3t_{s}(52+9t_{s}(-6+t_{s}(-8+t_{s}(287+9t_{s}(-74+35t_{s})))))\big{)}{}_{2}F_{1}{\Big{(}\frac{1}{3},\frac{8}{3},\frac{10}{3},\frac{2t_{s}}{-1+t_{s}}\Big{)}}$
$+$ $(1+t_{s})\big{(}130+9(-1+t_{s})t_{s}(-26+3t_{s}(-10-6t_{s}+45t_{s}^{2}))\big{)}{}_{2}F_{1}{\Big{(}\frac{4}{3},\frac{8}{3},\frac{10}{3},\frac{2t_{s}}{-1+t_{s}}\Big{)}}\big{\}}\bigg{\}}\Biggr{]}$
So, we can rewrite the above components follow the power laws $\sigma_{xy}^{(B^{2},\Omega)}(\omega)\propto\frac{\alpha_{J}^{\frac{2}{J}}}{\mu^{\frac{2}{J}}\hbar^{2(1-\frac{1}{J})}}$.
Refer to caption  Refer to caption  Refer to caption
Figure 6: The dependence of the optical conductivity on the tilt $t_{+}=0.5$ for the case of B $\perp t_{s}$. for (a) single WSM (b)double WSM and (c) triple WSM. The other parameters are the same as those of Fig.(1)
Refer to caption  Refer to caption  Refer to caption
Figure 7: The dependence of the optical conductivity on the tilt $t_{+}=0.5$ for the case of B $\perp t_{s}$. for (a) single WSM (b)double WSM and (c) triple WSM. The other parameters are the same as those of Fig.(1)
Refer to caption  Refer to caption  Refer to caption
Figure 8: The dependence of planar Hall conductivity on the tilt $t_{+}=0.5$ and $\gamma=\frac{\pi}{4}$ for (a) single WSM (b)double WSM and (c) triple WSM. The other parameters are the same as those of Fig.(1)
The magnetic orbital moment contributions of conductivities
for J=1
$\sigma_{xx}^{(B^{2},m)}(\omega)$ $=$ $\frac{\mathit{v}_{F}^{3}}{120\pi\mu^{2}}\sigma_{1}^{(B^{2})}\Bigl{[}(-3-6t_{s}^{2})\cos^{2}\gamma-\sin^{2}\gamma\Bigr{]}$
$\sigma_{zz}^{(B^{2},m)}(\omega)$ $=$ $\frac{\mathit{v}_{F}^{3}}{120\pi\mu^{2}}\sigma_{1}^{(B^{2})}(-1+t_{s}^{2})$ (76)
$\sigma_{xy}^{(B^{2},m)}(\omega)$ $=$ $\frac{\mathit{v}_{F}^{3}}{120\pi\mu^{2}}\sigma_{1}^{(B^{2})}(-2-6t_{s}^{2})\sin\gamma\cos\gamma$ (77)
for J=2
$\sigma_{xx}^{(B^{2},m)}(\omega)$ $=$ $\frac{\mathit{v}_{F}\alpha_{2}}{64\mu\hbar}\sigma_{1}^{(B^{2})}(2\cos^{2}\gamma-\sin^{2}\gamma)$ (78)
$\sigma_{zz}^{(B^{2},m)}(\omega)$ $=$ $\frac{\mathit{v}_{F}^{3}}{60\pi\mu^{2}}\sigma_{1}^{(B^{2})}(-1+t_{s}^{2})$ (79)
$\sigma_{xy}^{(B^{2},m)}(\omega)$ $=$ $\frac{\mathit{v}_{F}\alpha_{2}}{32\hbar\mu}\sin\gamma\cos\gamma\sigma_{1}^{(B^{2})}$ (80)
for J=3
$\begin{split}\sigma_{xx}^{(B^{2},m)}(\omega)&=\frac{\mathit{v_{F}}\alpha_{3}^{\frac{2}{3}}}{8\pi^{\frac{1}{2}}\hbar^{\frac{4}{3}}\mu^{\frac{2}{3}}}\sigma_{1}^{(B^{2})}\Biggl{[}-\frac{1}{24(1-t_{s})^{\frac{1}{3}}t_{s}^{5}\Gamma{(\frac{1}{6})}}\Gamma{(\frac{5}{3})}\bigg{\{}(1-t_{s})_{2}F_{1}{\Bigl{(}\frac{4}{3},\frac{5}{3},\frac{7}{3},\frac{2t_{s}}{-1+t_{s}}\Bigr{)}}\Big{(}6(-91+2t_{s}(-39+t_{s}(-13+\\ &3t_{s}(-5+2t_{s}(5+3t_{s})))))\cos^{2}\gamma+\big{(}-182+3t_{s}(-52-9(-3+t_{s})t_{s}(1+t_{s})))\big{)}\sin^{2}\gamma\Big{)}+_{2}F_{1}{\Bigl{(}\frac{1}{3},\frac{5}{3},\frac{7}{3},\frac{2t_{s}}{-1+t_{s}}\Bigr{)}}\\ &\big{\{}6(91+t_{s}(39+2t_{s}(-26+3t_{s}(-2+3t_{s}(-5+2(-2+t_{s})t_{s}))))\big{\}}\cos^{2}\gamma+(182+3t_{s}(26+t_{s}(-79\\ &+27t_{s}(-1+t_{s}+t_{s}^{2}))))\sin^{2}\gamma\big{\}}\bigg{\}}\Bigg{]}\end{split}$ (81)
$\sigma_{zz}^{(B^{2},m)}(\omega)=\frac{\mathit{v}_{F}^{3}\sigma_{1}^{(B^{2})}}{40\pi\mu^{2}}(-1+t_{s}^{2})\\$ (82)
$\begin{split}\sigma_{xy}^{(B^{2},m)}(\omega)&=\frac{\mathit{v}_{F}\alpha_{3}^{\frac{2}{3}}\sigma_{1}^{(B^{2})}}{8\pi^{\frac{3}{2}}\mu^{\frac{2}{3}}\hbar^{\frac{4}{3}}}{\sin\gamma\cos\gamma}\frac{\Gamma\big{(}\frac{-1}{3}\big{)}\Gamma\big{(}\frac{-1}{6}\big{)}}{1296(1-t_{s})^{\frac{1}{3}}t_{s}^{5}}\bigg{[}\big{\{}-364+3t_{s}\big{(}-52+t_{s}(25-3t_{s}(1+3t_{s}-(23+\\ &t_{s}(-19+8t_{s}))))\big{)}\big{\}}{{}_{2}F_{1}\big{(}\frac{1}{3},\frac{5}{3},\frac{7}{3},\frac{2t_{s}}{-1+t_{s}}\big{)}}-(1+t_{s})\big{\{}-384+3t_{s}\big{(}-104+t_{s}(-79+3t_{s}(-26+\\ &t_{s}(43+24t_{s})))\big{)}\big{\}}{{}_{2}F_{1}\big{(}\frac{4}{3},\frac{5}{3},\frac{7}{3},\frac{2t_{s}}{-1+t_{s}}\big{)}}\bigg{]}\end{split}$ (83)
So, we can rewrite the above components follow the power laws $\sigma_{zz}^{(B^{2},m)}(\omega)\propto\frac{J}{\mu^{2}}$ , $\sigma_{xx}^{(B^{2},m)}(\omega)\propto\frac{\alpha_{J}^{\frac{2}{J}}}{\mu^{\frac{2}{J}}\hbar^{2(1-\frac{1}{J})}}$ and $\sigma_{xy}^{(B^{2},m)}(\omega)\propto\frac{\alpha_{J}^{\frac{2}{J}}}{\mu^{\frac{2}{J}}\hbar^{2(1-\frac{1}{J})}}$.
Refer to caption  Refer to caption  Refer to caption
Figure 9: The frequency dependence of optical conductivity at tilt $t_{+}=0.5$, $\gamma=\frac{\pi}{4}$ and B = 1 T. for (a) single WSM (b)double WSM and (c) triple WSM.The other parameters are the same as those of Fig.(1)
It is noted that all the other magnetoconductivity components are zero. It is clear that the above conductivity equations are independent of chirality, i.e., the Weyl cones with opposite chiralities have the same contribution to the conductivity. The conductivity component $\sigma_{zz}^{(B^{2},m)}(\omega)$ is always negative and follow a power laws condition $\sigma_{zz}^{(B^{2},m)}(\omega)\propto J$, a result that will suppress the total conductivity compare to its Berry curvature parts.[see Fig(7)] .
In the present system, the planar Hall effect can take place [34, 35, 36, 33, 37, 38] and manifest itself in a nonzero conductivity $\sigma_{xy}^{(B^{2})}(\omega)$. Similar to the diagonal component of the conductivity, $\sigma_{xy}^{(B^{2})}(\omega)$ also consists of the contributions from the Berry curvature and the orbital magnetic moment. Figure (8) shows an effect of the tilt on the planar Hall magnetoconductivity. It is seen that the total planar Hall conductivity is suppressed in the case of single WSM when the orbit magnetic moment is present as shown in Fig.(8a). On the other hand, the total conductivity is enhanced with tilt parameter due to positive magnetic orbital moment for double and triple WSMs as shown in Fig.(8b) and Fig.(8c). Figure (9) shows the planar Hall magnetoconductivity as a function of the THz incident light at $\gamma=\pi/4$. The real and imaginary parts of total conductivities are suppressed in the case of single WSM due to negative contribution of real and imagnary parts of magnetic orbital moment conductivtes. For double and triple WSMs, the real and imagnary parts of conductivities are enhanced due to positive contributions of real and imaginary parts of orbital magnetic moment conductivites.
The $B^{2}$ dependence of the magnetoconductivity has been observed experimentally in the materials such as GdPtBi and TaP [37, 38]. TaS and GdPtBi have isotropic and quadratic band dispersions respectively. Therefore they are examples of single and double WSMs. Recently, it is also shown that in the materials GdPtBi, a very strong planar Hall effect has been reported, which is due to the Berry curvature and chiral anomaly contributions [37]. Besides chiral anomaly, the planar Hall effect may be induced by the orbital magnetic moment.
III.4 Hall conductivities $\sigma_{ab}^{(H,0)}$ and $\sigma_{ab}^{(H,B)}$
The intrinsic Hall effect is engendered by the Berry curvature, as presented by the second term of Eq. (17), for which further calculation gives
$\sigma_{ab}^{(H)}=\sigma_{ab}^{(H,0)}+\sigma_{ab}^{(H,B)}$ (84)
where
$\sigma_{ab}^{(H,0)}(\omega)$ $=$ $-\frac{e^{2}}{\hbar(2\pi)^{3}}\epsilon_{abc}\int d^{3}k\Omega_{c}^{s}f_{0}^{s}$ (85)
$\sigma_{ab}^{(H,0)}(\omega)$ $=$ $\frac{e^{2}}{\hbar(2\pi)^{3}}\epsilon_{abc}\int d^{3}k\Omega_{c}^{s}(\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B})\Bigl{(}\frac{\partial f_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}\Bigr{)}$ (86)
where $\epsilon_{abc}$ is the Levi-Civita symbol with $a,b,c,\in{x,y,z}$ The first term $\sigma_{ab}^{(H,0)}(\omega)$ in Eq. (84), referring to the anomalous Hall effect, is not equivalent to zero only in the system with broken time-reversal symmetry [27, 22]. While the second term $\sigma_{ab}^{(H,0)}(\omega)$stands for the ordinary Hall conductivity linear in B, which is the counterpart to a semiclassical description related to Landau level formation in the quantum limit [32].
Let the magnetic field in spherical polar co-ordinate system $\textbf{B}=(B_{x},B_{y},B_{z})$ with $B_{x}=B\sin\theta\cos\phi,B_{x}=B\sin\theta\sin\phi,B_{z}=B\cos\theta$. For a single Weyl node, the B-linear contribution to the Hall conductivity can be written as
$\sigma^{(H,B)}\\ =\left(\begin{array}[]{ccc}0&\sigma_{1}^{H}\cos\theta&-\sigma_{1}^{H}\sin\theta\sin\phi\\ -\sigma_{1}^{H}\cos\theta&0&\sigma_{1}^{H}\sin\theta\cos\phi\\ \sigma_{1}^{H}\sin\theta\sin\phi&-\sigma_{1}^{H}\sin\theta\cos\phi&0\end{array}\right)$ (87)
where
for J=1
$\sigma_{1}^{H}=-\frac{e^{3}}{\hbar}\frac{B\mathit{v}_{F}}{24\pi^{2}\mu}$ (88)
for J=2
$\sigma_{1}^{H}=\frac{e^{2}}{\hbar}\frac{Be}{4\pi}\frac{(-2+t_{s}^{2}+2\sqrt{1-t_{s}^{2}})}{t_{s}^{4}\mathit{v}_{F}\hbar}\alpha_{2}$ (89)
for J=3
$\begin{split}\sigma_{1}^{H}=&-\frac{{e^{3}}B\alpha_{3}^{\frac{2}{3}}\mu^{\frac{1}{3}}}{\hbar^{\frac{7}{3}}{4\pi^{\frac{3}{2}}}\mathit{v}_{F}}\frac{(3)^{\frac{1}{2}}\Gamma{(\frac{5}{6})}}{(1-t_{s})^{\frac{4}{3}}t_{s}^{3}\Gamma{(\frac{4}{3})}}\Bigg{[}(-1+t_{s})(-35+t_{s}(5+3t_{s})){}_{2}F_{1}{\bigg{(}\frac{-2}{3},\frac{5}{3},\frac{7}{3},\frac{2t_{s}}{-1+t_{s}}\bigg{)}}+\\ &\big{(}-35+t_{s}(-10+t_{s}(8+t_{s}))\big{)}{}_{2}F_{1}{\bigg{(}\frac{1}{3},\frac{5}{3},\frac{7}{3},\frac{2t_{s}}{-1+t_{s}}\bigg{)}}\Bigg{]}\end{split}$ (90)
So, we can rewrite the above components follow the power laws $\sigma_{1}^{H}\propto\frac{\alpha_{J}^{\frac{2}{J}}\mu^{(1-\frac{2}{J})}}{\hbar^{2(1-\frac{1}{J})}}$.
In contrast to the single WSM node, the B-linear contribution to the Hall conductivity of multi-WSMs depends on the tilt parameter. The linear Hall conductivity is independent of $t_{s}$ in the case of a single WSM while in the case double and triple cases, its depend on $t_{s}$ as shown in Fig.(10).
Refer to caption  Refer to caption  Refer to caption
Figure 10: The dependence of the optical conductivity on the tilt $t_{+}=0.5$ for the case of B $\parallel t_{s}$. for (a)single WSM (b) double WSM and (c) triple WSM. The other parameters are the same as those of Fig.(1)
IV Second order non-linear response of multi-WSMs
Now, we explore the second-order nonlinear magneto-optical response of Weyl semimetals. Substituting Eq.(9) into Eq.(10), neglecting the Lorentz force term, and retaining terms up to second order in E, we obtain
$\frac{1}{\hbar D}[-e\textbf{E}-\frac{e^{2}}{\hbar}(\textbf{E}\cdot\textbf{B})\Omega_{\bm{k}}^{s}]\cdot\frac{\partial\tilde{f}_{1}^{s}}{\partial\bm{k}}-i\omega\tilde{f}_{2}^{s}=-\frac{\tilde{f}_{2}^{s}}{\tau}$ (91)
Solve for $\tilde{f}_{2}^{s}$, we obtain
$\tilde{f}_{2}^{s}=\frac{\tau}{(1-2i\omega\tau)}\frac{1}{\hbar D}[e\textbf{E}+\frac{e^{2}}{\hbar}(\textbf{E}\cdot\textbf{B})\Omega_{\bm{k}}^{s}]\cdot\frac{\partial\tilde{f}_{1}^{s}}{\partial\bm{k}}$ (92)
Substituting Eq.(15) into Eq.(92) and retan terms up to first order in B, we obtain
$\tilde{f}_{2}^{s}$ $=$ $\frac{e\tau^{2}}{\hbar(1-i\omega\tau)(1-2i\omega\tau)}\biggl{\{}\textbf{E}\cdot\frac{\partial}{\partial\bm{k}}\biggr{[}\Bigl{(}e\textbf{E}+\frac{e^{2}}{\hbar}(\textbf{E}\cdot\textbf{B})\Omega_{\bm{k}}^{s}-\frac{e^{2}}{\hbar}(\textbf{B}\cdot\Omega_{\bm{k}}^{s})\textbf{E}\Bigr{)}\cdot\bm{\mathit{v}}_{\bm{k}}^{s}\frac{\partial f_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}-\frac{e}{\hbar}\textbf{E}\cdot\frac{\partial}{\partial\bm{k}}\biggl{(}\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B}\frac{\partial f_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}\biggr{)}\biggr{]}$ (93)
$+$ $\biggl{[}\frac{e^{2}}{\hbar}(\textbf{E}\cdot\textbf{B})\Omega_{\bm{k}}^{s}-\frac{e^{2}}{\hbar}(\textbf{B}\cdot\Omega_{\bm{k}}^{s})\textbf{E}\biggr{]}\cdot\frac{\partial}{\partial\bm{k}}\biggl{(}\textbf{E}\cdot\bm{\mathit{v}}_{\bm{k}}^{s}\frac{\partial f_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}\biggr{)}\biggr{\}}$
Now the electric current density at the frequency $2\omega$ is given by
$\bm{j}_{1}$ $=$ $-\frac{e}{(2\pi)^{3}}\int d^{3}k\Bigl{[}\bm{\mathit{\tilde{v}}}_{\bm{k}}^{s}+\frac{e}{\hbar}(\Omega_{\bm{k}}^{s}\cdot\bm{\mathit{\tilde{v}}}_{\bm{k}}^{s})\textbf{B}\Bigr{]}\tilde{f}_{2}^{s}$ (94)
- $\frac{e^{2}}{2\pi)^{3}\hbar}\int d^{3}k\textbf{E}\times\Omega_{\bm{k}}^{s}\tilde{f}_{1}^{s}$
According to the definition of second harmonic conductivity, this equation should be written in the form
$\textbf{j}(2\omega)=\sigma(2\omega)\textbf{E}(\omega)\textbf{E}(\omega)$ (95)
where $\sigma(2\omega)$ is the second harmonic conductivity.
IV.1 Second harmonic conductivity $\sigma_{abc}^{0}$ in the absence of magnetic field
In this subsection, we calculate the second harmonic current of the Weyl semimetals system in the absence of magnetic fields $\textbf{B}=0$. Inserting Eq. (93) into the first term of Eq. (94), the second harmonic conductivity tensor can be written as
$\sigma_{abc}^{0}(2\omega)=\frac{-\tau^{2}}{(1-i\omega\tau)(1-2i\omega\tau)}\frac{e^{3}}{\hbar(2\pi)^{3}}\int d^{3}k\frac{\partial\mathit{v}_{a}^{s}}{\partial k_{c}}\mathit{v}_{b}^{s}\Bigl{(}-\frac{\partial f_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}\Bigr{)}$ (96)
Except for the aaz, aza, and zaa (a = x, y, z) components of the second harmonic conductivity tensor are nonzero and all other components equal to zero [22, 39]
for J=1
$\sigma_{zzz}^{0}(2\omega)$ $=$ $\frac{\mu}{\pi\hbar^{2}}\sigma_{DL}\Bigl{[}\frac{-6+4t_{s}^{2}}{t_{s}^{3}}-\frac{3(1-t_{s}^{2})}{t_{s}^{4}}\ln\frac{1-t_{s}}{1+t_{s}}\Bigr{]}$
$\sigma_{xxz}^{0}(2\omega)$ $=$ $\frac{\mu}{2\pi\hbar^{2}}\sigma_{DL}\Bigl{[}\frac{6}{t_{s}^{3}}+\frac{3-t_{s}^{2}}{t_{s}^{4}}\ln\frac{1-t_{s}}{1+t_{s}}\Bigr{]}$ (98)
for J=2
$\sigma_{zzz}^{0}(2\omega)$ $=$ $\frac{\mathit{v}_{F}^{2}}{4\hbar\alpha_{2}}\sigma_{DL}\Bigl{[}\frac{2-2\sqrt{1-t_{s}^{2}}}{t_{s}^{3}}+\frac{-3+2\sqrt{1-t_{s}^{2}}}{t_{s}}\Bigr{]}$
$\sigma_{xxz}^{0}(2\omega)$ $=$ $\frac{\mu}{\pi\hbar^{2}}\sigma_{DL}\Bigl{[}\frac{6}{t_{s}^{3}}+\frac{3-t_{s}^{2}}{t_{s}^{4}}\ln\frac{1-t_{s}}{1+t_{s}}\Bigr{]}$ (100)
for J=3
$\sigma_{zzz}^{0}(2\omega)$ $=$ $\frac{\mathit{v}_{F}^{2}\sigma_{DL}}{2\pi^{\frac{1}{2}}\hbar^{\frac{2}{3}}\alpha_{3}^{\frac{2}{3}}\mu^{\frac{1}{3}}}\frac{-2\Gamma\big{(}\frac{4}{3}\big{)}}{3(1-t_{s})^{\frac{2}{3}}\Gamma{(\frac{11}{6})}}\bigg{[}(-1+t_{s}){}_{2}F_{1}{\Big{(}\frac{2}{3},\frac{4}{3},\frac{8}{3},\frac{2t_{s}}{-1+t_{s}}\Big{)}}+{}_{2}F_{1}{\Big{(}\frac{2}{3},\frac{7}{3},\frac{11}{3},\frac{2t_{s}}{-1+t_{s}}\Big{)}}\bigg{]}$ (101)
$\sigma_{xxz}^{0}(2\omega)$ $=$ $\frac{\mu}{2\pi\hbar^{2}}\sigma_{DL}\Bigl{[}\frac{18}{t_{s}^{3}}-\frac{3(-3+t_{s}^{2})}{t_{s}^{4}}\ln\Bigl{[}\frac{1-t_{s}}{1+t_{s}}\Bigr{]}\Bigr{]}$ (102)
where $\sigma_{DL}=\frac{e^{3}\tau^{2}}{(1-2i\omega\tau)(1-i\omega\tau)4\pi\hbar}$ is the Drude-like frequency dependent complex conductivity.
So, we can rewrite the above components follow the power laws $\sigma_{zzz}^{(0)}(2\omega)\propto\frac{\mu^{-(1-2/J)}}{\alpha_{J}^{\frac{2}{J}}\hbar^{\frac{2}{J}}}$ and $\sigma_{xxz}^{(0)}(2\omega)\propto\frac{J\mu}{\hbar^{2}}$.
Refer to caption  Refer to caption  Refer to caption
Figure 11: The nonlinear conductivities for the process of second harmonic generation of the tilt $t_{+}=0.5$, for (a) single WSM (b)double WSM and (c) triple WSM. The other parameters are the same as those of Fig.(1)
The second harmonic conductivity tensor satisfies the relation $\sigma_{zxx}^{0}(2\omega)=\sigma_{zyy}^{0}(2\omega)=\sigma_{xzx}^{0}(2\omega)=\sigma_{yzy}^{0}(2\omega)=\sigma_{xxz}^{0}(2\omega)=\sigma_{yyz}^{0}(2\omega)$, and it does not depend on the chirality of Weyl node. The total second harmonic conductivity in tilted Weyl semimetals is the sum of a pair of Weyl nodes. For the case with tilt inversion symmetry $t_{+}=-t_{-},\sigma_{abc}(2\omega)=0$. For the case with broken tilt inversion symmetry $t_{+}=t_{-},\sigma_{abc}(2\omega)\neq 0$.
It is noted to see from Fig.(11) that the $\sigma_{abc}^{0}(2\omega)$ becomes exactly zero when $t_{s}\rightarrow 0$. The presence of the finite tilt is needed to get the second harmonic generation in mWSMs. It is clear from Fig.(11) that $\sigma_{zzz}^{0}(2\omega)$ is more sensitive to tilt that $\sigma_{xxz}^{0}(2\omega)$. Fig.(13) shows the frquency variation of multi-WSMs.
IV.2 Second harmonic conductivity $\sigma_{abc}^{B}$ in the presence of linear magnetic field
Inserting Eq. (93) into the first term of Eq. (94), we write the linear B dependent second harmonic conductivity tensor
$\sigma_{abc}^{B}(2\omega)$ $=$ $\frac{\tau^{2}}{(1-i\omega\tau)(1-2i\omega\tau)}\frac{e^{3}}{\hbar(2\pi)^{3}}\int d^{3}k\biggl{\{}\frac{\partial\mathit{v}_{a}^{s}}{\partial k_{c}}\Bigl{[}-\frac{eB_{b}}{\hbar}(\Omega_{\bm{k}}^{s}\cdot\bm{\mathit{v}}_{\bm{k}}^{s})+\frac{e\mathit{v}_{b}^{s}}{\hbar}(\Omega_{\bm{k}}^{s}\cdot\bm{B})\Bigr{]}-\frac{eB_{c}}{\hbar}\frac{\partial}{\partial\bm{k}}\cdot(\mathit{v}_{a}^{s}\Omega_{\bm{k}}^{s})\mathit{v}_{b}^{s}$ (103)
$+$ $\frac{e}{\hbar}\frac{\partial[\mathit{v}_{a}^{s}(\Omega_{\bm{k}}^{s}\cdot\bm{B})]}{\partial k_{c}}\mathit{v}_{b}^{s}-\frac{eB_{a}}{\hbar}\frac{\partial(\Omega_{\bm{k}}^{s}\cdot\bm{\mathit{v}}_{\bm{k}}^{s})}{\partial k_{c}}\mathit{v}_{b}^{s}+\frac{\partial^{2}(\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B})}{\hbar\partial k_{a}\partial k_{c}}\mathit{v}_{b}^{s}-\frac{\partial^{2}\mathit{v}_{a}^{s}}{\hbar\partial k_{a}\partial k_{c}}(\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B})\biggr{\}}\Bigl{(}-\frac{\partial f_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}\Bigr{)}$
We will explore Eq.(103) with following two cases:
Refer to caption  Refer to caption  Refer to caption
Figure 12: The frequency dependence of optical conductivity at tilt $t_{+}=0.5$, for (a) single WSM (b)double WSM and (c) triple WSM. The other parameters are the same as those of Fig.(1)
Case-I In the case of $\textbf{B}\parallel t_{s}$ , for a single Weyl node, we obtain the magnetoconductivity components
for J=1
$\sigma_{xzx}^{B}(2\omega)$ $=$ $\frac{\mathit{v}_{F}^{2}}{12\pi^{2}\mu}s\sigma_{2}^{(B)}$ (104)
$\sigma_{zxx}^{B}(2\omega)$ $=$ $-\frac{\mathit{v}_{F}^{2}}{12\pi^{2}\mu}s\sigma_{2}^{(B)}$ (105)
for J=2
$\sigma_{xzx}^{B}(2\omega)$ $=$ $\frac{\alpha_{2}}{2\pi}s\sigma_{2}^{(B)}\frac{\sqrt{1-t_{s}^{2}}}{t_{s}^{4}}[2-t_{s}^{2}-2\sqrt{1-t_{s}^{2}}]$
$\sigma_{zxx}^{B}(2\omega)$ $=$ $-\frac{\alpha_{2}}{2\pi}s\sigma_{2}^{(B)}\frac{1}{t_{s}^{4}}[2-t_{s}^{2}-2\sqrt{1-t_{s}^{2}}]$ (107)
for J= 3
$\sigma_{xzx}^{B}(2\omega)=\frac{2s\sigma_{2}^{(B)}\alpha_{3}^{\frac{2}{3}}\mu^{\frac{1}{3}}}{\hbar^{\frac{4}{3}}\pi^{\frac{1}{2}}}\frac{1}{{3(3)^{\frac{1}{2}}(1-t_{s})^{\frac{4}{3}}t_{s}^{3}{\big{(}\frac{1+t_{s}}{1-t_{s}}\big{)}^{\frac{2}{3}}}\Gamma{(\frac{1}{6})}\Gamma{(\frac{1}{3})}}}\Bigg{[}(-1+t_{s})\bigg{\{}-70+3t_{s}\big{\{}-10+t_{s}\big{(}11\\ +18t_{s}(-1+\bigg{(}\frac{1+t_{s}}{1-t_{s}}\bigg{)}^{\frac{2}{3}}\bigg{(}-1+\frac{2}{1+t_{s}}\bigg{)}^{\frac{2}{3}})\big{)}\big{\}}\bigg{\}}+7(-10+9t_{s}^{2}){{}_{2}F_{1}{\bigg{(}\frac{2}{3},1,\frac{7}{3},\frac{2t_{s}}{-1+t_{s}}\bigg{)}}}\Bigg{]}$ (108)
$\begin{split}\sigma_{zxx}^{B}(2\omega)=&\frac{-6s\sigma_{2}^{(B)}\alpha_{3}^{\frac{2}{3}}\mu^{\frac{1}{3}}}{\hbar^{\frac{4}{3}}\pi^{\frac{1}{2}}}\frac{1}{{81(3)^{\frac{1}{2}}(1-t_{s})^{\frac{4}{3}}t_{s}^{5}(\frac{1+t_{s}}{1-t_{s}})\Gamma{(\frac{1}{6})}\Gamma{(\frac{4}{3})}}}\\ &\Bigg{[}{(-1+t_{s})\bigg{\{}2912+3t_{s}\big{\{}416+3t_{s}\big{(}-293+t_{s}(-89+31t_{s})\big{)}\big{\}}\bigg{\}}+\bigg{(}2912-3885t_{s}^{2}+1080t_{s}^{4}\bigg{)}{{}_{2}F_{1}{\bigg{(}\frac{2}{3},1,\frac{7}{3},\frac{2t_{s}}{-1+t_{s}}}}\bigg{)}}\Bigg{]}\end{split}$ (109)
where $\sigma_{2}^{(B)}=\frac{e^{4}\tau^{2}B}{\hbar^{2}(1-2i\omega\tau)(1-i\omega\tau)}$
So, we can rewrite the above components follow the power laws $\sigma_{xzx}^{(B)}(2\omega)\propto\frac{\alpha_{J}^{\frac{2}{J}}\mu^{(1-\frac{2}{J})}}{\hbar^{2(1-\frac{1}{J})}}$ and $\sigma_{zxx}^{(B)}(2\omega)\propto\frac{\alpha_{J}^{\frac{2}{J}}\mu^{(1-\frac{2}{J})}}{\hbar^{2(1-\frac{1}{J})}}$.
We can check that other nonzero second harmonic conductivity components satisfy the relations: $\sigma_{zxx}^{B}(2\omega)=\sigma_{zyy}^{B}(2\omega)=\sigma_{xzx}^{B}(2\omega)=\sigma_{yzy}^{B}(2\omega)$. In contrast to single WSM, the B-linear contribution to the second harmonic conductivity depends on the tilt. Summing the conductivity over the Weyl cones with opposite chirality cancels each other, leading to the disappearance of the total B-linear contribution to the second harmonic conductivity.
Case-II In the case of $\textbf{B}\perp t_{s}$ , we get the nonzero conductivity components
for J=1
$\sigma_{zxz}^{(B)}(2\omega)$ $=$ $s\frac{\mathit{v}_{F}^{2}}{12\pi^{2}\mu}\sigma_{2}^{(B)}\cos\gamma$ (110)
$\sigma_{xzz}^{(B)}(2\omega)$ $=$ $-s\frac{\mathit{v}_{F}^{2}}{12\pi^{2}\mu}\sigma_{2}^{(B)}\cos\gamma$ (111)
$\sigma_{zyz}^{(B)}(2\omega)$ $=$ $s\frac{\mathit{v}_{F}^{2}}{12\pi^{2}\mu}\sigma_{2}^{(B)}\sin\gamma$ (112)
$\sigma_{yzz}^{(B)}(2\omega)$ $=$ $-s\frac{\mathit{v}_{F}^{2}}{12\pi^{2}\mu}\sigma_{2}^{(B)}\sin\gamma$ (113)
for J=2
$\sigma_{zxz}^{(B)}(2\omega)$ $=$ $s\frac{\mathit{v}_{F}^{2}}{6\pi^{2}\mu}\sigma_{2}^{(B)}\cos\gamma$ (114)
$\sigma_{xzz}^{(B)}(2\omega)$ $=$ $-s\frac{2\mathit{v}_{F}^{2}}{3\pi^{2}\mu}\sigma_{2}^{(B)}\cos\gamma$ (115)
$\sigma_{zyz}^{(B)}(2\omega)$ $=$ $s\frac{\mathit{v}_{F}^{2}}{6\pi^{2}\mu}\sigma_{2}^{(B)}\sin\gamma$ (116)
$\sigma_{yzz}^{(B)}(2\omega)$ $=$ $-s\frac{2\mathit{v}_{F}^{2}}{3\pi^{2}\mu}\sigma_{2}^{(B)}\sin\gamma$ (117)
for J=3
$\sigma_{zxz}^{(B)}(2\omega)$ $=$ $s\frac{\mathit{v}_{F}^{2}}{4\pi^{2}\mu}\sigma_{2}^{(B)}\cos\gamma$ (118)
$\sigma_{xzz}^{(B)}(2\omega)$ $=$ $-s\frac{\mathit{v}_{F}^{2}}{4\pi^{2}\mu}\sigma_{2}^{(B)}\cos\gamma$ (119)
$\sigma_{zyz}^{(B)}(2\omega)$ $=$ $s\frac{\mathit{v}_{F}^{2}}{4\pi^{2}\mu}\sigma_{2}^{(B)}\sin\gamma$ (120)
$\sigma_{xzz}^{(B)}(2\omega)$ $=$ $-s\frac{\mathit{v}_{F}^{2}}{4\pi^{2}\mu}\sigma_{2}^{(B)}\sin\gamma$ (121)
So, we can rewrite the above components follow the power laws $\sigma_{zxz}^{(B)}(2\omega)\propto\frac{J}{\mu}$ , $\sigma_{xzz}^{(B)}(2\omega)\propto\frac{J}{\mu}$ , $\sigma_{zyz}^{(B)}(2\omega)\propto\frac{J}{\mu}$ and $\sigma_{yzz}^{(B)}(2\omega)\propto\frac{J}{\mu}$.
In this case, the B-linear contribution to the second harmonic conductivity is dependent on the chirality but independent of the tilt. Summing the conductivity over the Weyl cones with opposite chirality cancels each other, leading to the disappearance of the total B-linear contribution to the second harmonic conductivity.
IV.3 The second order nonlinear Hall conductivity $\sigma_{abc}^{(H,0)}$ in the absence of magnetic field
In this subsection, we study the second-order nonlinear Hall effect of Weyl semimetals without magnetic field. Inserting Eq. (16) into the second term in Eq. (94) and taking B = 0, we obtain the nonlinear Hall conductivity.
$\sigma_{abc}^{(H,0)}=\epsilon_{acd}\frac{e^{3}\tau}{\hbar^{2}(1-i\omega\tau)}D_{bd}$ (122)
where $D_{bd}$ is called Berry dipole. It is a first-order moment of the Berry curvature[40].
One can calculate the non-zero components of Berry curvature dipole
$D_{bd}=\frac{\hbar}{(2\pi)^{3}}\int d^{3}k\Omega_{d}^{s}\mathit{v}_{b}^{s}\Bigl{(}-\frac{\partial f_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}\Bigr{)}$ (123)
for J=1
$D_{xx}=D_{yy}$ $=$ $-s\frac{1}{16\pi^{2}}\Bigl{[}\frac{2}{t_{s}}+\frac{1-t_{s}^{2}}{t_{s}^{3}}\ln\frac{1-t_{s}}{1+t_{s}}\Bigr{]}$ (124)
$D_{zz}$ $=$ $-s\frac{1}{8\pi^{2}}\frac{t_{s}^{2}-1}{t_{s}^{3}}\Bigl{[}2t_{s}+\ln\frac{1-t_{s}}{1+t_{s}}\Bigr{]}$ (125)
for J=2
$D_{xx}=D_{yy}$ $=$ $-s\frac{1}{8\pi^{2}}\Bigl{[}\frac{2}{t_{s}^{2}}+\frac{1-t_{s}^{2}}{t_{s}^{3}}\ln\frac{1-t_{s}}{1+t_{s}}\Bigr{]}$ (126)
$D_{zz}$ $=$ $-s\frac{1}{4\pi^{2}}\frac{t_{s}^{2}-1}{t_{s}^{3}}\Bigl{[}2t_{s}+\ln\frac{1-t_{s}}{1+t_{s}}\Bigr{]}$ (127)
for J= 3
$D_{xx}=D_{yy}$ $=$ $-s\frac{1}{8\pi^{2}}\Bigl{[}\frac{3}{t_{s}^{2}}+\frac{3(1-t_{s}^{2})}{2t_{s}^{3}}\ln\frac{1-t_{s}}{1+t_{s}}\Bigr{]}$ (128)
$D_{zz}$ $=$ $-s\frac{1}{4\pi^{2}}\frac{t_{s}^{2}-1}{t_{s}^{3}}\Bigl{[}3t_{s}+\frac{3}{2}\ln\frac{1-t_{s}}{1+t_{s}}\Bigr{]}$ (129)
In the case of DWSMs and Triple-WSMs, the magnitude of Berry curvature dipole components is twice and thrice respectively the values of single WSMs and it has the same dependence on the chirality, as well as on the tilt. Therefore, contributions from a pair of Weyl nodes with opposite chirality exactly cancel each other [22, 40, 41].
IV.4 Linear B-contribution to the second order nonlinear Hall conductivity $\sigma_{abc}^{(H,B)}$
Now, we focus on the second-order nonlinear Hall effect in a weak magnetic field. Inserting Eq. (16) into the second term of Eq. (94) and retaining terms up to the first power of B, one obtains complex nonlinear Hall conductivity
$\sigma_{abc}^{(H,B)}=\epsilon_{acd}\frac{e^{3}\tau}{\hbar^{2}(1-i\omega\tau)}\bigl{[}D_{bd}^{\Omega}+D_{bd}^{m}\bigr{]}$ (130)
where
$D_{bd}^{\Omega}=\frac{e}{(2\pi)^{3}}\int d^{3}k\Omega_{d}^{s}[B_{b}(\Omega_{\bm{k}}^{s}\cdot\bm{\mathit{v}}_{\bm{k}}^{s})-\mathit{v}_{b}^{s}(\Omega_{\bm{k}}^{s}\cdot\bm{B})]\Bigl{(}-\frac{\partial f_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}\Bigr{)}$ (131)
$D_{bd}^{m}=\frac{1}{(2\pi)^{3}}\int d^{3}k\frac{\partial\Omega_{d}^{s}}{\partial k_{b}}(\bm{m}_{\bm{k}}^{s}\cdot\bm{B})\Bigl{(}-\frac{\partial f_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}\Bigr{)}$ (132)
where $D_{bd}^{\Omega}$ and $D_{bd}^{m}$ are Berry curvature dipole contributions due to the Berry curvature and the orbital magnetic moment, respectively.
Again we will consider the following two cases.
Case I In the case of $\textbf{B}\parallel t_{s}$, one obtains the Berry curvature dipole components.
for J=1
$D_{xx}^{\Omega}=D_{yy}^{\Omega}$ $=$ $-\frac{t_{s}}{120}D_{2}^{(B)}$ (133)
$D_{zz}^{m}$ $=$ $\frac{t_{s}}{60}D_{2}^{(B)}$ (134)
$D_{xx}^{m}=D_{yy}^{m}$ $=$ $\frac{t_{s}}{60}D_{2}^{(B)}$ (135)
$D_{zz}^{m}$ $=$ $-\frac{t_{s}}{30}D_{2}^{(B)}$ (136)
for J=2
$D_{xx}^{\Omega}=D_{yy}^{\Omega}=D_{zz}^{\Omega}=0$
$D_{xx}^{m}=D_{yy}^{m}=D_{zz}^{m}=0$ (138)
In this case, all components are zero, which is finite in the case of isotropic WSM[22].
for J=3
$D_{xx}^{\Omega}=D_{yy}^{\Omega}=\frac{D_{2}^{(B)}}{8}\frac{\alpha_{3}^{\frac{2}{3}}\pi^{3/2}\mu^{\frac{4}{3}}}{\hbar^{\frac{4}{3}}\mathit{v_{F}^{2}}}\Bigl{[}\frac{\pi^{\frac{5}{2}}2766400(\frac{1}{1-t_{s}})^{\frac{1}{3}}((-1+t_{s}){{}_{2}F_{1}{(\frac{-2}{3},\frac{8}{3},\frac{16}{3},\frac{2t_{s}}{-1+t_{s}})}}+{{}_{2}F_{1}{(\frac{1}{3},\frac{8}{3},\frac{16}{3},\frac{2t_{s}}{-1+t_{s}})}})}{729(3)^{\frac{1}{2}}t_{s}\Gamma{(\frac{25}{6})}\Gamma{(\frac{19}{3})}}\Bigr{]}$ (139)
$D_{zz}^{\Omega}=-2\frac{D_{2}^{(B)}}{8}\frac{\alpha_{3}^{\frac{2}{3}}\pi^{3/2}\mu^{\frac{4}{3}}}{\hbar^{\frac{4}{3}}\mathit{v_{F}^{2}}}\Bigl{[}\frac{\pi^{\frac{5}{2}}2766400(\frac{1}{1-t_{s}})^{\frac{1}{3}}((-1+t_{s}){{}_{2}F_{1}{(\frac{-2}{3},\frac{8}{3},\frac{16}{3},\frac{2t_{s}}{-1+t_{s}})}}+{{}_{2}F_{1}{(\frac{1}{3},\frac{8}{3},\frac{16}{3},\frac{2t_{s}}{-1+t_{s}})}})}{729(3)^{\frac{1}{2}}t_{s}\Gamma{(\frac{25}{6})}\Gamma{(\frac{19}{3})}}\Bigr{]}$ (140)
$D_{xx}^{m}=D_{yy}^{m}=\frac{D_{2}^{(B)}}{8}\frac{\alpha_{3}^{\frac{2}{3}}\pi^{3/2}\mu^{\frac{4}{3}}}{\hbar^{\frac{4}{3}}\mathit{v_{F}^{2}}}\frac{4(2)^{\frac{1}{3}}\pi}{729(1-t_{s})^{\frac{1}{3}}\Gamma{\bigl{(}\frac{7}{3}\bigr{)}}}\Bigl{[}6561\Gamma{\bigl{(}\frac{2}{3}\bigr{)}^{2}{{}_{2}F_{1}{\Bigl{(}\frac{1}{3},\frac{5}{7},\frac{7}{3},\frac{2t_{s}}{-1+t_{s}}\Bigr{)}}}}+\\ \frac{2(3)^{\frac{1}{2}}\pi\Gamma{\bigl{(}\frac{-1}{3}\bigr{)}}\bigg{\{}(65+3t_{s}(26+3t_{s}(1+72t_{s}))){{}_{2}F_{1}{\Bigl{(}\frac{1}{3},\frac{8}{3},\frac{10}{3},\frac{2t_{s}}{-1+t_{s}}\Bigr{)}}}-(1+t_{s})(65+9t_{s}(13+12t_{s})){{}_{2}F_{1}{\Bigl{(}\frac{4}{3},\frac{8}{3},\frac{10}{3},\frac{2t_{s}}{1+t_{s}}\Bigr{)}}}\bigg{\}}}{t_{s}^{3}\Gamma{\bigl{(}\frac{10}{3}\bigr{)}}}\Bigr{]}$ (141)
$D_{zz}^{m}=-\frac{D_{2}^{(B)}}{8}\frac{\alpha_{3}^{\frac{2}{3}}\pi^{3/2}\mu^{\frac{4}{3}}}{\hbar^{\frac{4}{3}}\mathit{v_{F}^{2}}}\frac{4\pi^{2}(2)^{\frac{1}{3}}\Gamma{(\frac{2}{3})}}{81(3)^{\frac{1}{2}}(1-t_{s})^{\frac{1}{3}}t_{s}^{4}\Gamma{(\frac{7}{2})^{2}}}\Bigl{[}(-91+3t_{s}(-13+t_{s}(29+9t_{s}))){{}_{2}F_{1}{\Bigl{(}\frac{1}{3},\frac{5}{3},\frac{7}{3},\frac{2t_{s}}{-1+t_{s}}\Bigr{)}}}\\ -(1+t_{s})(-91-78t_{s}+9t_{s}^{2}){{}_{2}F_{1}{\Bigl{(}\frac{4}{3},\frac{5}{3},\frac{7}{3},\frac{2t_{s}}{-1+t_{s}}\Bigr{)}}}\Bigr{]}$ (142)
where, $D_{2}^{(B)}=\frac{eB\hbar\mathit{v}_{F}^{2}}{\pi^{2}\mu^{2}}$.
Refer to caption  Refer to caption
Figure 13: The nonlinear Hall conductivities for the process of second harmonic generation as a function of the incident photon frequency at tilt $t_{+}=0.5$ , B=1 T and $\gamma=\frac{\pi}{4}$ for (a) single WSM (b)triple WSM.. The other parameters are the same as those of Fig.(1)
Case-II In the case of $\textbf{B}\perp t_{s}$ , we get the nonzero components
for J=1
$D_{zx}^{\Omega}$ $=$ $-\frac{t_{s}}{20}D_{2}^{(B)}\cos\gamma$ (143)
$D_{xz}^{\Omega}$ $=$ $\frac{3t_{s}}{40}D_{2}^{(B)}\cos\gamma$ (144)
$D_{zy}^{\Omega}$ $=$ $-\frac{t_{s}}{20}D_{2}^{(B)}\sin\gamma$ (145)
$D_{yz}^{\Omega}$ $=$ $\frac{3t_{s}}{40}D_{2}^{(B)}\sin\gamma$ (146)
and
$D_{zx}^{m}$ $=$ $-\frac{t_{s}}{40}D_{2}^{(B)}\cos\gamma$ (147)
$D_{zy}^{m}$ $=$ $-\frac{t_{s}}{40}D_{2}^{(B)}\sin\gamma$ (148)
for J= 2
$D_{zx}^{\Omega}$ $=$ $-2\frac{t_{s}}{20}D_{2}^{(B)}\cos\gamma$ (149)
$D_{xz}^{\Omega}$ $=$ $\frac{\pi\mu\alpha_{2}}{16\hbar v_{F}^{2}}t_{s}D_{2}^{(B)}\cos\gamma$ (150)
$D_{zy}^{\Omega}$ $=$ $-2\frac{t_{s}}{20}D_{2}^{(B)}\sin\gamma$ (151)
$D_{yz}^{\Omega}$ $=$ $\frac{\pi\mu\alpha_{2}}{16\hbar v_{F}^{2}}t_{s}D_{2}^{(B)}\sin\gamma$ (152)
and
$D_{zx}^{m}=-2\frac{t_{s}}{40}D_{2}^{(B)}\cos\gamma$ (153)
$D_{zy}^{m}=-2\frac{t_{s}}{40}D_{2}^{(B)}\sin\gamma$ (154)
for J=3
$D_{zx}^{\Omega}$ $=$ $-\frac{3t_{s}}{20}D_{2}^{(B)}\cos\gamma$ (155)
$D_{zy}^{\Omega}$ $=$ $-\frac{3t_{s}}{20}D_{2}^{(B)}\sin\gamma$ (156)
$D_{zx}^{m}$ $=$ $-\frac{3t_{s}}{40}D_{2}^{(B)}\cos\gamma$ (157)
$D_{zy}^{m}$ $=$ $-\frac{3t_{s}}{40}D_{2}^{(B)}\sin\gamma$ (158)
Refer to caption  Refer to caption
Figure 14: The angle dependence of the nonlinear Hall conductivities at $\omega=5THz$ for (a) single WSM (b) triple WSM.. The other parameters are the same as those of Fig.(1)
$\begin{split}D_{xz}^{\Omega}&=\frac{D_{2}^{(B)}}{8}\frac{\alpha_{3}^{\frac{2}{3}}\pi^{3/2}\mu^{\frac{4}{3}}}{\hbar^{\frac{4}{3}}\mathit{v_{F}^{2}}}\frac{\Gamma\big{(}\frac{-1}{3}\big{)}^{2}}{486(2)^{\frac{2}{3}}t_{s}^{4}(1-t_{s})^{\frac{1}{3}}\Gamma\big{(}\frac{7}{3}\big{)}}\Bigg{[}\big{\{}-91+3t_{s}(-13+t_{s}(43+3t_{s}(5+2t_{s}(1+9t_{s}))))\big{\}}{{}_{2}F_{1}\big{(}\frac{1}{3},\frac{5}{3},\frac{7}{3},\frac{2t_{s}}{-1+t_{s}}\big{)}}\\ &-(1+t_{s})\big{\{}91+3t_{s}\big{(}-26+t_{s}(17+6t_{s}(2+3t_{s}))\big{)}\big{\}}{{}_{2}F_{1}\big{(}\frac{4}{3},\frac{5}{3},\frac{7}{3},\frac{2t_{s}}{-1+t_{s}}\big{)}}\Bigg{]}\cos\gamma\end{split}$ (159)
$\begin{split}D_{yz}^{\Omega}=\frac{D_{2}^{(B)}}{8}\frac{\alpha_{3}^{\frac{2}{3}}\pi^{3/2}\mu^{\frac{4}{3}}}{\hbar^{\frac{4}{3}}\mathit{v_{F}^{2}}}\Bigg{[}\frac{(3)^{\frac{1}{2}}\Gamma{(\frac{2}{3})^{3}}\big{\{}(-13+15t^{2}+54t^{4}){{}_{2}F_{1}{(\frac{1}{3},\frac{5}{3},\frac{10}{3},\frac{2t}{-1+t_{s}})}}-(1+t_{s}){{}_{2}F_{1}{(\frac{4}{3},\frac{5}{3},\frac{10}{3},\frac{2t_{s}}{-1+t})}}\big{\}}}{{14(2)^{\frac{2}{3}}t_{s}^{3}(1-t_{s})^{\frac{1}{3}}}}\Bigg{]}\sin\gamma\end{split}$ (160)
So, we can rewrite the above components follow the power laws $D_{zx}^{\Omega}\propto J$ , $D_{xz}^{\Omega}\propto\frac{\alpha_{J}^{\frac{2}{3}}\mu^{(2-\frac{2}{J})}}{\hbar^{(2-\frac{2}{J})}}$ , $D_{zy}^{\Omega}\propto J$ and $D_{yz}^{\Omega}\propto\frac{\alpha_{J}^{\frac{2}{3}}\mu^{(2-\frac{2}{J})}}{\hbar^{(2-\frac{2}{J})}}$.
One can check that $D_{zx}^{m}=D_{xz}^{m}$ and $D_{zy}^{m}=D_{yz}^{m}$ and rest of the components will vanish. The contributions of all components are renormalized due to parameters $\alpha_{2}$ and $\mathit{v}_{F}$ of DWSMs as compared to single WSMs [22].
The nonlinear Hall effect can be modulated by the polarization of the incident light, as discussed in Ref.[22]. Using Eq. (130), the electric current is rewritten in the form of
$\textbf{j}(2\omega)=\frac{e^{3}\tau}{\hbar^{2}(1-i\omega\tau)}(\hat{D}\cdot\textbf{E})\times\textbf{E}$ (161)
Assume that an electromagnetic wave propagates in the x direction:
$\textbf{E}(\textbf{r},t)=|\textbf{E}(\omega)|Re[|\psi\rangle e^{i(qx-\omega t)}$ (162)
where
$|\psi\rangle=\Bigl{(}\begin{array}[]{c}\psi_{y}\\ \psi_{z}\end{array}\Bigr{)}=\Bigl{(}\begin{array}[]{c}\sin\theta e^{i\alpha_{y}}\\ \cos\theta e^{i\alpha_{z}}\end{array}\Bigr{)}$ (163)
is the Jones vector in the y-z plane with phases $\alpha_{y},\alpha_{z}$ , and the amplitudes $E_{y}=|\textbf{E}|\sin\theta$ and $E_{z}=|\textbf{E}|\cos\theta$ . Inserting Eq. (162) into Eq. (161), we obtain the nonlinear Hall current
$j_{x}(2\omega)=\frac{e^{3}\tau}{\hbar^{2}(1-i\omega\tau)}\frac{D_{yy}^{(B)}-D_{zz}^{(B)}}{2}\sin 2\theta e^{(\alpha_{y}+\alpha_{z})}|\textbf{E}|^{2}$ (164)
Following the same procedure of Ref.([22]), Eq.(164) can be rewritten as $j_{x}\sim(D_{yy}-D_{zz})E_{y}E_{z}$ . Hence conductivity will be
for J=1
$\sigma_{x}^{(H,B)}(2\omega)=\frac{e^{3}\tau}{(1-i\omega\tau)80\hbar^{2}}{D_{2}^{(B)}t_{s}\sin 2\gamma}$ (165)
for J=2
$\sigma_{x}^{(H,B)}(2\omega)=0$ (166)
for J=3
$\begin{split}\sigma_{x}^{(H,B)}(2\omega)&={\frac{e^{3}\tau}{\hbar^{2}(1-i\omega\tau)}}\frac{D_{2}^{(B)}}{8}\frac{\alpha_{3}^{\frac{2}{3}}\pi^{3/2}\mu^{\frac{4}{3}}}{\hbar^{\frac{4}{3}}\mathit{v_{F}^{2}}}\sin 2\gamma\biggl{[}-\frac{1}{24(3\pi)^{\frac{1}{2}}t_{s}(-1+t_{s}^{2})\Gamma{(\frac{7}{3})}\Gamma{(\frac{25}{6})}}19(1-t_{s})^{\frac{1}{3}}\Bigg{\{}5(-1+t_{s})\biggl{(}3\Bigl{(}\frac{1}{1-t_{s}}\Bigr{)}^{\frac{2}{3}}\\ &+(1-t_{s})^{\frac{1}{3}}+(1-t_{s})^{\frac{1}{3}}\biggr{)}{}_{2}F_{1}{(\frac{-2}{3},\frac{8}{3},\frac{16}{3},\frac{2t_{s}}{-1+t_{s}})}\Bigl{\{}5\biggl{(}3\Bigl{(}\frac{1}{1-t_{s}}\Bigr{)}^{\frac{2}{3}}+(1-t_{s})^{\frac{1}{3}}\biggr{)}\\ &+5(1-t_{s})^{\frac{1}{3}}t_{s}-\Bigl{(}\frac{1}{1-t_{s}}\Bigr{)}^{\frac{2}{3}}t_{s}^{2}\Bigr{\}}{}_{2}F_{1}{(\frac{1}{3},\frac{8}{3},\frac{16}{3},\frac{2t_{s}}{-1+t_{s}})}\Bigg{\}}\biggr{]}\end{split}$ (167)
Fig.(13) shows the variation of linear-B second harmonic conductivities $\sigma_{x}^{(H,B)}(2\omega)$ as function of incdent photon frequency $\omega$. In contrast to single and triple WSMs, this component has zero value for double WSM. In contrast to the case of $B=0$, the B-linear contribution to the nonlinear Hall conductivity is independent of the chirality, and the odd function of $t_{s}$ . Only in the system with broken tilt inversion symmetry ($t_{+}=t_{-}$ ), the conductivity $\sigma_{x}^{(H,B)}(2\omega)\neq 0$. From Eq. (92), evidently, $\sigma_{x}^{(H,B)}(2\omega)$ reaches its maximum when the polarization direction $\theta=\pm\pi/4$ and vanishes at $\theta=0,\pi/2$, as further reflected in Fig.(14).
Recent experiments shows the second harmonic optical response generates gaint second-order nonlinear optical polarizability, $\chi^{2}$, in transition metal monopnictides (TMMPs) such as TaAs, a class of noncentrosymmetric materials [42]. Another experment reports large non-linear Hall effect(NLHE) due to gigantic Berry curvature dipole density as generated by tilted Weyl cones near the Fermi level in a model ferroelectric Weyl semimetal in doped $(Pb_{1-x}Sn_{x})_{1-y}In_{y}Te$. The effective Berry curvature dipole derived from the experimentally observed nonlinear Hall voltage follows a scaling law with carrier concentration, which is consistent with the simplest form of the Berry curvature dipole expected for the Weyl monopoles [43].
V Conclusion
We have calculated the magneto-optical conductivities of gapless type-I tilted-multi Weyl semimetals in the presence of orbital magnetic moment in linear and non-linear responses within the semiclassical Boltzmann approach and compared it with the case of single WSMs. We have found that conductivity components are renormalized in a non-trivial manner as power laws. Our calculations would be tested in mWSMs materials.
Appendix A DETAILS OF THE CALCULATIONS USING SPHERICAL POLAR COORDINATES
In this paper, we focus on the n-doped multi-Weyl semimetals with a positive chemical potential $\mu$. In general, one can decompose the momentum k into parallel and perpendicular parts:
$k_{x}$ $=$ $\Bigl{(}\frac{k\sin\theta}{\alpha_{J}}\Bigr{)}^{\frac{2}{J}}\cos\phi$ (168)
$k_{x}$ $=$ $\Bigl{(}\frac{k\sin\theta}{\alpha_{J}}\Bigr{)}^{\frac{2}{J}}\sin\phi$ (169)
$k_{z}$ $=$ $\frac{k}{v_{F}}\cos\theta$ (170)
The Jacobian of the transformation is $\mathcal{J}=\frac{1}{Jv_{F}\sin\theta}\left(\frac{k\sin\theta}{\alpha_{J}}\right)^{2/J}$, which has been used for analytical expressions of conductivity elements of multi-WSMs.